A Complete Bibliography of ACM Transactions on Management Information Systems (TMIS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 June 2024 Version 1.42

Title word cross-reference

#swineflu [KSS14].

H [HC11]. K [LKZ^{+}15]. l [PCK21].

-diversity [PCK21]. -means [LKZ^{+}15].

1 [CSSD22]. 10 [HL11]. 10-K [HL11]. 19 [ABE^{+}21, BEK^{+}23, BMR^{+}21, GBK23, JNDZ23, KJM23, LKPZ23, LZLL21, PPG^{+}23, SJ23, WZZ22, ZJT^{+}23, ZBX^{+}21, ZKMK21].

2015 [BJJK17]. 2017 [JRYY19]. 2018 [DFJ20].

4.0 [ALS22, CTED22].

6G [CFH22].

Absorptive [HL21]. Academic [AMZ20]. Access [RSM^{+}17, RSM^{+}21, WHEC22, XWLC19]. accumulation [NM12]. Accuracy [KM23]. Accuracy-Interpretability [KM23]. Accurate [EV13]. ACM [Che10]. Across [KJM23]. Actionable [MOR22]. Activities [CRW19, WZZS20]. Activity [PCK21, ZBX^{+}21, AC10]. Acute [MH13]. Ad [MSD17, SWB19]. Adapt [EV13]. Adapter [SK12]. Adapting [ZZX^{+}18]. Adaptive [JBEM15, WWL^{+}22]. add [DDGR11]. Address [AY22]. Adoption [KS13, WHEC22, ZTD^{+}13]. Adversarial [CLX^{+}23, SMY20, ZELC22]. Adversarial-Based [CLX^{+}23]. Adverse [CJH22]. Advertising [LZX16, LCC17]. Advice [RB23]. Advisory [NMS12, SRS11]. Affecting [WHEC22]. Affective [KC13]. Affordance [WHEC22]. Affordances [ZJT^{+}23]. Age [PHL^{+}21]. agencies [SRS11]. Agency [BR15]. Agent [FBP24, LAYR13, YHLW15, GIYZ11]. Agent-Based [FBP24]. Agent-Oriented [YHLW15]. Aggregation [PHL^{+}21]. agile [CRAH10]. AI [CFH22, DFL^{+}23, PGPB19, ZZYT21]. AI-augmented [DFL^{+}23]. AI-Based [CFH22, PGPB19]. Aid [ZJT^{+}23]. AIDCOV [ZKMK21]. Alerts [SMY20]. Algorithm [LZZ^{+}22, LZX^{+}22b, MSD17, WEM^{+}13, XWC21, HL11]. Algorithmic [AY22]. Algorithms [BDLC20, EV13, GUH16, LCS^{+}20, SKC21]. Alignment [CNJ13, UL13, WWL^{+}24, XWC21]. Alliance [BMS17]. Allocation [CWY^{+}23, CN23, HDZ^{+}23, NHL22]. Allocations [FBP24]. Amateur [XZBZ21]. Ambiguity [DSZ17, LHT15]. Among [BMRW15, SZX^{+}23, SKT11]. Analyses [ZCL15]. Analysis [AMZ20, BSPP18, BMV21, BR15, CXG^{+}22, DXC23, KZ20, KOH^{+}21, KJRD16, KUH21, NST21, PWS^{+}15, WR12, YZJ^{+}19, ZGG13, ZAZC18, AC10, KY11]. Analytical [WWL^{+}15, WZZ22]. Analytics [BMS17, CGS12, CSSD22, LCC13, LSSD22, LLF^{+}22, QGRT21, RB23, SKC21, SJC15, SJ23]. Analytics-Based [RB23]. Analyze [LZLL21]. Analyzing [HC11, JHC19, NYS12, SCR23]. Android [PMH^{+}20, STBI22]. Annealing [MHT19, NFVY^{+}22]. Annealing-based [MHT19]. Annotation [CLX^{+}23]. Anomaly [NSK^{+}21, RASHD22]. Anonymization [PCK21]. Anonymous [WCX21]. Answering [KF19]. ante [LSS20]. App [MSD17]. Application [KS13, RCW22, ZNJ19, PZH11]. Applications [HSK17, LCS^{+}20, ZTD^{+}13, ZCL15]. Appraisal [KM23]. Approach [ARLEG17, BRH^{+}15, CNJ13, CAL22, Ema19, FKK^{+}23, HLYW22, HBK^{+}20, MRKE23, NST21, PZW24, QGRT21, QvdHT24, STASH20, SJC15, SJ23, SSY16, TZ20, WR12, XZM^{+}22]. Approaches [CFH22, HZZ12]. Apriori [MSD17]. archetypes [DW11]. Architecture [JFI^{+}23, LZX^{+}22b]. Areas [LCWH24]. Argumentation [RB23]. Art [ZAZC18]. Artful [DM15]. article [LR11]. Artifacts [SSY16, BFV12]. Artificial [HS23, LZX^{+}22b, SKC20, SKC21, WRZ22, ZKMK21]. Aspect [DXC23, TZ20]. Aspect-based [DXC23]. Assembly [CCWL22]. Assessing [OUAC21, Sie23]. Assessment [FKBW15, KZ19]. Assignment [RSM^{+}17]. assignments [GIYZ11]. Assistive [EV13]. Assurance [FKBW15]. Asynchronous [SK12]. Attack [COS^{+}21]. Attacks [KUH21, WXR10]. Attention [CXG^{+}22, XLLX18, HLYW22]. Attribute [RSM^{+}21]. Attribute-Based [RSM^{+}21]. Attribution [STASH20]. Auctions [SWB19]. Audit [LYC16]. Auditable [CMM^{+}22]. Auditing [MRKE23]. Audits [ZML^{+}13]. Augmentation [SNS^{+}20]. augmented [DFL^{+}23]. Australian [KC13, PWS^{+}15]. Authentication [PGS22]. Author [STASH20]. Autocorrelations [ZTD^{+}13]. Autoencoder [CLX^{+}23, WZZ22]. Automated [ACC^{+}21, COS^{+}21, LHT15, NSK^{+}21, AHS^{+}23]. Automatically [HYYY22]. Automating [MOR22]. Availability [XLZ^{+}16]. Available [YYJ14]. Aware [LT16, UTL20, ALS22, QGRT21]. Awareness [HBK^{+}20]. away [BKSZ10]. AZEmo [YZJ^{+}19].

Balancing [ALS22]. Ballistics [AVBC23]. Bank [LCWH24]. Based [CLX^{+}23, CFH22, CCWL22, CXG^{+}22, DBC^{+}14, DMJ^{+}13, DTL^{+}22, FBP24, FKBW15, GYZY22, JRYY19, KTF15, KZ20, KKB^{+}22, LHT15, LKZ^{+}15, LYC16, LSM^{+}21, LZX^{+}22a, MSB20, MSD17, PGPB19, PGS22, PZW24, RCW22, RSM^{+}17, RSM^{+}21, RB23, UTL20, WH13, XLLX18, XZM^{+}22, ZELC22, ZLY^{+}15, AMZ20, CJH22, CMM^{+}22, DXC23, Ema19, JFI^{+}23, KM23, LZZ^{+}22, MCCC22, MAKH^{+}11, MHT19, RASHD22, RW11, WCX21, WCL^{+}21, ZLC12]. Basket [ULST23]. Batch [POtH22]. Batch-Processing [POtH22]. Bayesian [BSA14, Orm13, SZX^{+}23]. Before [SJ23]. Behavior [DMJ^{+}13, LKZ^{+}15, LAYR13, SZX^{+}23, Uhl11, WXR10]. Behavior-Based [LKZ^{+}15]. Behavioral [AY22, ARLEG17, LZX16, RADS12]. Behaviors [ALJH12]. Benchmark [ZAZC18]. Benefits [MH13]. Between [Kak17, LC19, PZW24]. Bias [AY22]. Bidders [SWB19]. Big [KKB^{+}22, LST22, LCF^{+}22, LLF^{+}22, TSCT18]. Binary [LCS^{+}20, ZTD^{+}13]. Biometric [SG21]. Blacklisting [HBK^{+}20]. Blackouts [PPC^{+}20]. Bladder [FKK^{+}23]. Blockchain [AVBC23, CMM^{+}22, JFI^{+}23, MRKE23, WCX21]. Blockchain-based [JFI^{+}23, WCX21]. Blockchains [MWV^{+}18]. Blood [HLYW22]. Booking [HLCY22]. Bookmarking [WR12]. Books [ZZ23]. Botnet [LKZ^{+}15]. Boundary [JCMR15]. BPM [SSY16]. Branches [LCWH24]. Breaking [AFD^{+}21]. Bride [SCT^{+}13]. Bridge [PZW24]. Broadband [PCLL19]. broadcasting [RW11]. broader [NB11]. Budget [CWY^{+}23]. Building [Goo14, HBL14]. Business [ALGA24, BRH^{+}15, BSPP18, CGS12, CNJ13, DFL^{+}23, FKBW15, GUH16, GSV15, JJ19, LKPZ23, LCC13, MWV^{+}18, MH13, OAHY21, SSY16, SJKP17, UL13, XLZ^{+}16, YHLW15, RW11, SRS11]. Business-to-Consumer [SJKP17].

CAE [Goo14]. Can [LAYR13].

Candlestick [TQ14]. Capacity [HL21]. capital [DFP11]. Capping [HSK17]. CAPTCHA [ZELC22]. Care [KC13, MH13, ZML^{+}13]. Case [AVBC23, BTZ17, BMFL15, NM15, WCL^{+}21]. Cashless [Sre20]. Casual [NMS12]. Catastrophic [PHL^{+}21]. Catching [HYG19]. Categories [KZ20]. Categorization [TLK20]. Cautionary [PHL^{+}21]. centered [PZH11]. Centric [AY22]. Certification [SJKP17]. Chain [KTF15]. Challenges [BDLC20, EFPS^{+}22, HS23, KCSTM18, MWV^{+}18, BGMS11, Che11]. Change [ARLEG17, LAYR13, LTC^{+}12]. Changes [SJKP17]. Channel [JFI^{+}23]. characteristics [AZ12]. Characterization [PMH^{+}20]. Charging [PGS22]. Charts [TQ14]. Chat [DMJ^{+}13, STASH20, GIYZ11]. Chat-Based [DMJ^{+}13]. Checking [JADT15]. Chest [JNDZ23, ZKMK21]. Chinese [YZJ^{+}19]. Chip [LZX^{+}22b]. Choice [HZZ12]. Choreography [CMM^{+}22]. Choreography-based [CMM^{+}22]. CIOs [DW11]. CircleCI [RCM^{+}22]. Cities [LCF^{+}22]. Clarity [TW17]. Classes [LS17]. Classification [CLX^{+}23, DSZ17, KTF15, LGW^{+}22, GIYZ11, HL11]. Classifier [BEK^{+}23]. Classifiers [BSA14]. Classifying [STBI22]. Click [GSS16]. Click-Through-Rate [GSS16]. Climbing [NFVY^{+}22]. Clinical [BMFL15, MO13, PWS^{+}15, QGRT21, SYS^{+}13]. Cloud [DTL^{+}22, EFK^{+}20, MAKH^{+}11, MRKE23, NHL22, OAHY21, XWLC19]. Cloud-based [MAKH^{+}11]. Clustering [LHT15]. CNN [JNDZ23]. Co [RASHD22, SXH^{+}24, WWL^{+}24]. Co-evolution-based [RASHD22].

Co-Medication [SXH^{+}24]. Co-occurrence [WWL^{+}24]. Coherence [LT16]. Coherence-Aware [LT16]. Collaboration [ZJT^{+}23, Kan11, LR11]. Collaborative [HLYW22, PZH11]. Collection [CWY^{+}23]. Collective [LZK21]. Colleges [Goo14]. Combination [BSA14]. Combining [LZX16]. Commerce [TLK20, YZJ^{+}19]. Communication [DMJ^{+}13, KC13, KSS14, SK12, VSRU13]. Communities [LSS20, WWL^{+}15, WZZS20]. community [SKT11]. Compact [CLYH19, XWC21]. Company [AGL23]. Comparative [ETHJ22, PWS^{+}15]. Comparison [ZGG13]. Competitive [LZX^{+}22a, ZGG13]. complementary [DFP11]. Complexity [COS^{+}21, JL15, KS15, KUH21]. Compliance [JADT15, MCCC22, NM15, OUAC21]. Complications [KS15]. components [MCTC11]. Composite [ULST23]. Composition [QvdHT24, SK12, ZLY^{+}15, ZNJ19]. Compression [LYC16]. Computation [GSS16]. Computational [DBC^{+}14, LCF^{+}22]. Computer [DFP11]. Computers [CN23]. Computing [DTL^{+}22, EFK^{+}20, HDZ^{+}23, WZW22, XWLC19]. Concentration [GSS16]. Concept [LHT15, vdLSFEA23]. Concept-Based [LHT15]. Conceptual [LS17, PBT18]. Concerns [LZLL21]. Conditions [CJH22]. Conducting [SKC21]. Confidentiality [EFPS^{+}22]. Configurable [POtH22]. Confucianism [HBL14]. Congestion [WCL^{+}21]. Congruence [CSS17]. Connecting [SKC21]. Connections [SZX^{+}23]. Consortium [WCX21]. Constrained [RSM^{+}17]. Constraint [HLCY22]. Constraints [RSM^{+}15, XLZ^{+}16]. Consultations [vdLSFEA23]. Consumer [MPBH20, SJKP17, YYJ14, Uhl11, ZLC12]. containing [ARV11]. Contemporary [DJS18]. Content [BMV21, LT16, WR12, YYJ14]. Context [LST22, LL24, PFD^{+}13, UTL20]. Context-Aware [UTL20]. Context-Sensitive [PFD^{+}13]. Contexts [Kak17]. Contextual [LZX16, SNS^{+}20]. Contingency [SKT15]. Continuous [TWC16]. Continuous-Space [TWC16]. Contractual [NM15]. Contrasting [ZTD^{+}13]. Contributed [YYJ14]. Contributions [Goo14]. contributors [SKT11]. Control [AHS^{+}23, DM15, HBL14, OUAC21, RSM^{+}17, RSM^{+}21]. Controlling [HSK17]. Convolutional [WCL^{+}21]. Cooperative [RASHD22]. coordinate [MM12]. Coordination [SKT15]. Core [CCWL22]. corporate [AC10]. Correlated [CAL22, PGPB19]. CoRSAI [ABE^{+}21]. Counter [DZP^{+}24]. Counteracting [ZELC22]. County [ZBX^{+}21]. County-Level [ZBX^{+}21]. COVID [ABE^{+}21, BEK^{+}23, BMR^{+}21, GBK23, JNDZ23, KJM23, LSM^{+}21, LZLL21, PPG^{+}23, SJ23, WZZ22, ZJT^{+}23, ZBX^{+}21, ZKMK21, LKPZ23]. COVID-19 [ABE^{+}21, BEK^{+}23, BMR^{+}21, GBK23, JNDZ23, KJM23, LZLL21, PPG^{+}23, SJ23, WZZ22, ZJT^{+}23, ZBX^{+}21, ZKMK21, LKPZ23]. COVID-Safe [LSM^{+}21]. COVIDAL [BEK^{+}23]. Created [LS17]. Creation [SK12]. Credential [PGS22]. Credit [LTC^{+}12, Sie23]. Crime [AVBC23]. Criteria [ZNJ19]. Cross [NJC^{+}22]. Cross-Modality [NJC^{+}22]. Cruising [HLCY22]. CT [ABE^{+}21, JNDZ23]. Cues [DMJ^{+}13]. Culture [HBL14]. Customer [LP16, YRUP22, PHCS11]. Customers [PZW24, ULST23, GIYZ11]. Customization [KF19]. Cyber [GYZY22, Goo14, HBK^{+}20, KZ20, LCF^{+}22, PGPB19, PPC^{+}20, PHL^{+}21, SMY20, ZELC22]. Cyber-Insured [PGPB19]. Cyber-physical [LCF^{+}22].

Cyber-Physical-Social [GYZY22]. Cyber-Risk [PHL^{+}21]. Cybercrime [PCLL19]. Cybersecurity [AMZ20, AHS^{+}23, OUAC21, RASHD22, SKC20]. Cyberworld [SCT^{+}13].

Daily [PCK21]. Dark [ZELC22]. Data [BRH^{+}15, BMFL15, CWY^{+}23, DTL^{+}22, ETHJ22, GYZY22, HLG14, LZK21, LTL23, LST22, LT16, LCF^{+}22, LLF^{+}22, MO13, MSU^{+}20, MOR22, NSK^{+}21, PGPB19, PCK21, PMH^{+}20, QvdHT24, RCW22, SKC21, SXH^{+}24, SCR23, TSCT18, WHEC22, WRZ22, ZZYT21, AZ12, FCKG10, MAKH^{+}11]. Data-Driven [BRH^{+}15, PMH^{+}20]. Data-oriented [ETHJ22]. Databases [KUH21]. Dataflow [GSV15]. Dataset [KKB^{+}22]. Datasets [RASHD22, STBI22]. DAtt [CXG^{+}22]. Deal [BMRW15]. Deceit [EDBK21]. Decentralized [MRKE23, PGS22]. Deceptive [DMJ^{+}13]. Deciphering [ZLC12]. Decision [ARV11, CXG^{+}22, CN23, CSSD22, HLCY22, KOH^{+}21, LKPZ23, LSSD22, MO13, vdLSFEA23]. Decision-Support [vdLSFEA23]. Decisions [HCLN12, KS13, YRUP14]. Declarative [DM15]. Decomposition [DLN^{+}21]. Deduplication [MRKE23]. Deep [ABE^{+}21, CFH22, EDBK21, HLYW22, KM23, MSD23, SG21, TZ20, UTL20, WZZS20, WZZ22, XZM^{+}22, ZZ23]. Delegation [XWLC19]. Delivering [FN18]. Delivers [LAYR13]. Demand [YRUP22]. Denoising [CLX^{+}23]. Departments [WEM^{+}13]. Dependency [DSZ17, HMN^{+}16, NLLZ20]. Deploying [LCWH24]. Deployment [OAHY21]. Derivation [ZLY^{+}15]. Design [AVR21, ARLEG17, AHS^{+}23, DJS18, GH13, HH19, HS23, LTL23, LZX^{+}22b, LZX^{+}22a, NM12, OAHY21, PBT18, SB24, SRS11, Cha11, Che11]. Designing [RW11]. Detecting [AFD^{+}21, BTZ17, CRW19, DMJ^{+}13, STBI22]. Detection [BDLC20, COS^{+}21, CJH22, JNDZ23, KTF15, KKB^{+}22, LKZ^{+}15, MSB20, NSK^{+}21, PFD^{+}13, RASHD22, XLLX18, ZKMK21, LLK^{+}11, MAKH^{+}11]. Deter [ACC^{+}21, EYS^{+}17]. Determinants [CN23, GBK23]. Developer [KJRD16]. Developing [DAHND18, JHC19]. Development [BNSW15, CGS12, JCMR15, LCWH24, MFBK^{+}17, CRAH10]. Deviations [ZML^{+}13]. Devices [LZW^{+}19, MSB20]. DevOps [RCM^{+}22]. Diagnosis [BEK^{+}23, WZZ22, RADS12]. Dialogue [PZW24]. Differential [KLL19]. Difficulty [CWY^{+}23]. Digital [AVR21, AMZ20, BEK^{+}23, BGMS11, LCF^{+}22, SB24]. Dimensional [KKB^{+}22, CWY^{+}23, GYZY22]. Directions [LFS18, LCC13]. Disambiguation [Ema19]. Disciplinary [SKC20, SKC21]. Discipline [CGS12, SKC20, NM12]. disclosures [WXR10]. Discontinuity [TSCT18]. Discover [HYYY22]. Discovering [ARLEG17, TWC16, TSCT18]. Discovery [BSPP18, DM15, LZZ^{+}22, POtH22, RADS12, WR12]. Discrete [HZZ12]. Disease [SJ23, QGRT21]. Disentangling [ZJT^{+}23]. Dispatch [VSRU13]. Dispatch-Mediated [VSRU13]. Dispersed [SKT15]. Display [SWB19]. Disruptions [ZZYT21]. Dissonance [PZW24]. Distance [KKB^{+}22]. Distributed [MO13]. Distributions [FBP24]. diversity [PCK21]. DNN [MHC^{+}22]. Do [EYS^{+}17, HCLN12, DDGR11]. Doctor [vdLSFEA23]. Doctor-Patient [vdLSFEA23]. Document [LHT15]. Documents [ACC^{+}21, DZP^{+}24]. Does [WH13, LR11]. Dogs [Tuz11]. Domain [KF19, ULST23]. Don’t [ULST23]. Double [CXSW17, CXG^{+}22]. Double-Layer [CXSW17]. Driven [BRH^{+}15, CSSD22, GLC^{+}20, HZZ12, KKB^{+}22, LSSD22, LZX^{+}22b, LZX^{+}22a, PMH^{+}20]. Drivers [HLCY22, WXR10]. Drug [YYJ14]. DSSAE [WZZ22]. During [KOH^{+}21, SJ23, GBK23]. Dynamic [AGL23, CWY^{+}23]. Dynamics [BRH^{+}15, CNJ13, KS13, CRAH10].

E-Commerce [TLK20, YZJ^{+}19]. e-customers [GIYZ11]. Early [HL21, KSS14]. Economic [AY22, Sre20]. Economics [PHL^{+}21, PCLL19]. Ecosystem [BRH^{+}15]. Ecosystems [BSPP18]. ecoxight [BSPP18]. Edge [CN23, DTL^{+}22, MSB20]. Edge-Based [MSB20]. Edge-Cloud [DTL^{+}22]. Editorial [Che10, Che11, JL15]. Education [CGS12, HL21]. EEG [EDBK21]. Effect [LP16, OUAC21, ZB21, XZBZ21]. Effective [CLYH19, LZX16, LCC17, DW11]. Effects [CSS17, SXH^{+}24, Sre20, SCR23, YDS^{+}13]. Efficiency [COS^{+}21, SWB19]. Efficiency-Complexity [COS^{+}21]. Efficient [EV13, EDBK21, KF19, MRKE23, XWLC19]. Eggs [ULST23]. Ego [BR15]. EHR [MH13]. EHRs [vdLSFEA23]. Electric [PGS22]. Electrical [BTZ17]. Electricity [KCSTM18]. Electronic [ZML^{+}13]. Electronics [CCWL22]. Elicitation [Bur16, EFK^{+}20]. Eliciting [SKT11]. Embeddings [ACC^{+}21, ULST23]. Embodying [KC13]. Emergencies [ZJT^{+}23]. Emergency [BTZ17, VSRU13, WEM^{+}13]. Emerging [HYYY22, KCSTM18, BGMS11]. Emoticon [YZJ^{+}19]. Emotional [KC13]. Empathic [LAYR13]. Empirical [JCMR15, KTF15, KJRD16, MPBH20, WH13]. Employee [RSM^{+}17, RSM^{+}21]. Employees [OUAC21]. empowered [PGS22]. enabled [HDZ^{+}23, LKPZ23, SKC21]. Enabling [MRKE23, WR12]. Encounters [NMS12]. Energy [ALS22]. Energy-aware [ALS22]. Enforced [NMS12]. Engaging [AFD^{+}21]. Engine [LCC17, SYS^{+}13]. Engineering [CRD15, CMM^{+}22, JCMR15, JL15, NM15]. Engineers [Bur16]. Enhance [RB23]. Enhanced [EV13, NYS12]. Enhancement [BFV12]. Ensemble [STASH20]. Ensembles [STASH20]. Entropy [BMV21, WZZ22]. Entry [LZW^{+}19]. Environment [BR15, DBC^{+}14]. Environments [ALS22, DTL^{+}22, SB24]. Epidemiology [DBC^{+}14]. Equal [LS17]. Era [LKPZ23, PPG^{+}23]. ESG [AGL23]. Establishing [WRZ22]. Estate [KM23]. Estimates [PGPB19]. Estimating [LCWH24, SYS^{+}13]. Estimation [HSK17]. Evaluating [FN18]. Evaluation [BDLC20, MFBK^{+}17, ZAZC18, Cha11, MMW11]. Event [MHC^{+}22, MFBK^{+}17, PLW^{+}17, POtH22, RADS12]. Events [BTZ17, KTF15]. Evidence [SZX^{+}23, WH13]. Evolution [JL15, WZW22, RASHD22]. Evolutionary [ZNJ19]. Evolutive [LZZ^{+}22]. evolving [MAKH^{+}11]. Ex [LSS20]. Ex-ante [LSS20]. Ex-post [LSS20]. Examining [SJ23]. Exchange [BMV21, YDS^{+}13]. Exchanges [EYS^{+}17]. executive [MMW11]. Expectation [WEM^{+}13]. Explaining [Uhl11]. Exploiting [SZX^{+}23, TZLX21]. Exploration [BSPP18, DJS18, HL21]. Exploratory [BNSW15, Sre20]. Exploring [DLN^{+}21, HYYY22]. Expressways [WCL^{+}21]. Extracting [TZ20]. Extraction [BMV21, GYZY22].

Facebook [BMRW15]. Facilitating

[CRD15]. Facilities [KC13]. Factor [CN23]. Factorization [LCS^{+}20]. Factors [Sie23, WHEC22, HL11]. Fair [FBP24]. Fake [ACC^{+}21, DZP^{+}24]. Families [STBI22]. Family [JCMR15]. Fast [PFD^{+}13]. Feasibility [PPC^{+}20]. Feature [GYZY22, NSK^{+}21, RASHD22, SNS^{+}20]. Features [DSZ17, KZ19, LSM^{+}21]. Feedback [MHC^{+}22]. Fi [JFI^{+}23]. field [DGM^{+}10]. Finance [TWC16]. Financial [AVR21, BMV21, DXC23, LCWH24, LTC^{+}12, LP16, NMS12, RCFT18, TSCT18, ZZ23, LLKC11]. Findings [NMS12, MMW11]. Fine [HSK17]. Fine-Grained [HSK17]. Firm [ZZ23, ZCL15]. Firm-Related [ZCL15]. Firms [PGPB19]. First [PGPB19, Che10]. Fit [ZZX^{+}18]. Fitness [GH13]. Fitness-Utility [GH13]. Flexibility [HBL14]. Flexible [KOH^{+}21, XWLC19]. flower [Cha11]. Flows [DM15]. Flu [KSS14]. Forecasting [CXSW17, ZZ23]. Forensics [STASH20]. form [HL11]. Formality [XZBZ21]. Forums [BMR^{+}21, ZCL15]. forwards [DDGR11]. Foundations [LFS18, SB24]. Founded [FKBW15]. Four [PWS^{+}15]. Fourier [WZZ22]. Fractional [WZZ22]. Fracture [SXH^{+}24]. Framework [CXSW17, GVP21, HH19, HMN^{+}16, HLCY22, TW17, WWL^{+}15]. Frameworks [UTL20]. Free [LZW^{+}19]. Frequency [CXSW17, LGW^{+}22]. Frequent [CAL22, LZZ^{+}22]. Frontier [QvdHT24]. Fully [HCLN12]. Fusion [TZLX21]. Future [LFS18, LCWH24, ZB21, DGM^{+}10].

Gap [AMZ20, PZW24, WWL^{+}22]. Gathering [TW17]. GDPR [ZB21]. Generation [ACC^{+}21, LL24]. Generative [SMY20, ZELC22]. generic [SRS11]. Genetic [XWC21]. Geo [SNS^{+}20]. Geo-Contextual [SNS^{+}20]. German [BEK^{+}23]. Getting [CRD15]. give [BKSZ10]. Giving [RB23]. Globally [SKT15]. glyphs [Cha11]. Goals [YHLW15, SRS11]. goods [BKSZ10, BGMS11]. Grained [HSK17]. Grammars [LS17]. grand [Che11]. Graph [CXG^{+}22, Ema19, SJ23, WCL^{+}21]. Graph-based [Ema19]. Graphic [LCF^{+}22]. Grid [KCSTM18, DDGR11]. Group [BDD^{+}21, KY11]. Groupon [BMRW15]. Growth [Sre20, YDS^{+}13]. Guided [TZ20].

Gun [AVBC23].

HAIS [HS23]. Handed [HYG19, LZW^{+}19]. Handle [ZZYT21]. Health [ARLEG17, BMR^{+}21, EYS^{+}17, FN18, GBK23, LAYR13, WZZS20, YLA13, YYJ14, YDS^{+}13]. Health-Consumer-Contributed [YYJ14]. Healthcare [EFK^{+}20, LTL23, MSD23]. hedging [DDGR11]. Help [LAYR13]. Helpfulness [KZ19, NYS12]. Heterogeneous [ALS22, DTL^{+}22, HSK17, HDZ^{+}23, XZM^{+}22, GIYZ11]. Heuristics [ZNJ19]. Hierarchy [KZ19]. High [AFD^{+}21, CXSW17, CLYH19, DBC^{+}14, GYZY22, KKB^{+}22, NFVY^{+}22, ZBX^{+}21]. High-dimensional [GYZY22]. High-Engaging [AFD^{+}21]. High-Frequency [CXSW17]. High-Quality [CLYH19]. High-Resolution [ZBX^{+}21]. Higher [HL21]. Highly [MO13]. Hill [NFVY^{+}22]. Hip [SXH^{+}24]. History [XZM^{+}22]. Holes [BNSW15]. Home [HYG19]. Hong [LZLL21]. Hospital [SJ23]. Hospitals [BEK^{+}23, MH13, PWS^{+}15]. HOSVD [GYZY22]. Human [AY22, HS23, SCR23, DFP11]. Human-Artificial [HS23]. Human-Centric [AY22]. Human-in-the-Loop [SCR23]. Hybrid [QvdHT24].

IC [LZX^{+}22a]. ICT [KJM23]. Ideas

[DGM^{+}10]. Identification

[BNSW15, EDBK21, GVP21]. Identifying [BDD^{+}21, GHK^{+}20, LST22]. Identity [PGS22]. II [LSSD22]. IIOT [HDZ^{+}23]. Image [CLX^{+}23, KM23, NJC^{+}22]. Image-based [KM23]. Image-Text [NJC^{+}22]. Images [ZKMK21]. Imbalanced [MO13]. Impact [AZ12, BMR^{+}21, CNJ13, HBL14, LCC17, YRUP22, LR11]. Impacts [MH13, Che11, HC11]. Impediments [NM15]. imperfect [GIYZ11].

Implementation [CNJ13]. Implementing [FN18]. Improve [TW17]. Improved [EV13]. Improving [FKBW15, MSD17]. In-App [MSD17]. Incentive [MH13]. Incentives [SJKP17]. Incident [KZ20, WCL^{+}21]. Incident-related [WCL^{+}21]. Inclusion [AVR21]. Inclusive [AVR21]. Incorporating [DXC23]. Increasing [LS17]. Incremental [JHC19, LZZ^{+}22]. Indebted [LSS20]. Index [AVR21, HC11]. India [Sre20]. Individual [LZK21, ULST23]. Individual-Domain [ULST23]. Industrial [AHS^{+}23, COS^{+}21]. Industry [BTZ17, CCWL22, ALS22, CTED22]. INF [AVR21]. INF-PIE [AVR21]. Inference [Orm13, SYS^{+}13]. Influence [XLLX18]. Influence-Based [XLLX18]. Informatics [WHEC22]. Information [AHS^{+}23, BMV21, Bur16, CLYH19, CGS12, EYS^{+}17, FN18, GLC^{+}20, GBK23, HLCY22, HCLN12, JRYY19, KS13, KCSTM18, LLF^{+}22, NJC^{+}22, RCFT18, UL13, YDS^{+}13, HC11, Kan11, MMW11, NM12, PHCS11, WXR10, NB11]. InfoSec [MCCC22]. Infrared [FKK^{+}23]. Initial [FBP24]. Inner [SB24]. Innovation [GUH16, YRUP22]. Insecurity [MPBH20]. Insider [KUH21]. Insights [AY22, ZJT^{+}23]. inspired [DZP^{+}24]. instance [SG21]. Insurance [PPC^{+}20]. Insured [PGPB19]. Insurers [PHL^{+}21]. Integrating [AY22]. Integration [ALGA24, GSV15, GBK23, PBT18, RCM^{+}22]. Intellectual [ACC^{+}21]. Intelligence [AVBC23, CGS12, HS23, LCC13, LZX^{+}22b, LCF^{+}22, SKC20, SKC21, WHEC22, WRZ22, ZELC22, ZKMK21, RW11]. Intelligence-enabled [SKC21]. Intelligent [GLC^{+}20, HYYY22, LLF^{+}22]. Intention [CSS17]. Inter [ALGA24]. Inter-organizational [ALGA24]. Interaction [CRW19]. Interactions [JADT15, KJM23, KY11]. Interactive [BSPP18, DBC^{+}14, XWC21]. Interfirm [NST21]. Internet [CJH22, ETHJ22, MPBH20, STASH20]. Internet-scale [MPBH20]. Interpretability [KM23]. Interpretable [KM23, LCWH24, ZKMK21]. Interpretation [ABE^{+}21]. Interpretive [SCT^{+}13]. Interventions [LAYR13]. Introduction [BJJK17, CTED22, CSSD22, DFJ20, JRYY19, LTL23, LKPZ23, LSSD22, ZZYT21]. Intrusion [MSB20, SMY20]. Invested [LSS20]. Investigating [Kak17, SKT15, BKSZ10, LC19, LCC17]. Investigation [JCMR15, SCT^{+}13, MCTC11, WXR10]. Investment [KS13]. IoT [ALS22, COS^{+}21, CFH22, CTED22, DTL^{+}22, MSB20, PHL^{+}21]. IP [DZP^{+}24, SNS^{+}20]. Issue [BJJK17, CTED22, CSSD22, DFJ20, LTL23, LKPZ23, LSSD22, Che10]. issues [BGMS11]. IT-Business [CNJ13]. IT-enabled [LKPZ23]. IT-related [DFP11]. IT-Supported [NMS12]. Item [ZZA^{+}13]. Items [SYS^{+}13]. Itemset [NLLZ20]. Itemsets [NFVY^{+}22].

J [GYZY22]. J-HOSVD [GYZY22]. Job [SZX^{+}23, ZZX^{+}18]. Joint [LGW^{+}22, SZX^{+}23, ZZX^{+}18].

Key [WHEC22]. Keyboard [LZW^{+}19]. Keypoint [WWL^{+}24]. Keywords [TWC16]. Knowledge [DXC23, LSS20, LZZ^{+}22, LL24, QGRT21, Tuz11, WWL^{+}15, WH13, MM12, NM12, SKT11]. Knowledge-aware [QGRT21]. Knowledge-Sharing [WWL^{+}15]. Kong [LZLL21].

Label [CLX^{+}23]. labeling [HL11].

Language

[PBT18, TCS22, TWC16, XZBZ21, LLK^{+}11]. Large [FKBW15, STBI22]. Large-Scale [FKBW15, STBI22]. Latent [NST21, PZH11]. Launch [FBP24]. Layer [CXSW17]. Learning [ABE^{+}21, BEK^{+}23, BKMK23, BDD^{+}21, COS^{+}21, CFH22, DAHND18, EDBK21, FKK^{+}23, HLYW22, KF19, KM23, LLF^{+}22, MCCC22, MSD23, NJC^{+}22, QvdHT24, SZX^{+}23, SG21, TZ20, UTL20, WZZS20, XZM^{+}22, ZELC22, ZZX^{+}18]. Lending [Sie23]. Level [ZBX^{+}21]. Leveraging [LZK21]. Lightweight [JNDZ23]. Like [XZBZ21]. Likely [PPC^{+}20]. Likes [BMRW15, PHL^{+}21]. LiMS [JNDZ23]. LiMS-Net [JNDZ23]. linguistic [KY11]. Link [LFS18]. Links [LCC17]. Live [GIYZ11]. Live-chat [GIYZ11]. Load [ALS22]. Local [LCWH24, NLLZ20]. Local-Regional [LCWH24]. Locations [HLCY22, LZK21]. Lockdown [ZJT^{+}23]. Logs [PLW^{+}17, POtH22, TCS22]. Long [ZZ23]. Long-term [ZZ23]. Look [PGPB19]. Loop [SCR23]. Losses [KZ20]. Loyalty [Kak17].

Machine [BEK^{+}23, BDD^{+}21, COS^{+}21, DAHND18, FKK^{+}23, LLF^{+}22, MCCC22, MOR22, TLK20]. Machine-Actionable [MOR22]. Mail [SCT^{+}13]. Mail-Order-Bride [SCT^{+}13]. Majority [KLL19]. Making [CXG^{+}22, CSSD22, LKPZ23, LSSD22]. Malware [STBI22, MAKH^{+}11]. Management [AHS^{+}23, DFL^{+}23, ETHJ22, HMN^{+}16, LHT15, LKPZ23, LLF^{+}22, MWV^{+}18, MOR22, NJC^{+}22, OAHY21, Sie23, SSY16, WH13, XLZ^{+}16, Tuz11]. Manifesto [DFL^{+}23]. Manufacturers [PZW24]. Manufacturing [CCWL22, LZX^{+}22b, LZX^{+}22a, NHL22, RCW22]. Mapping [LHT15]. Market [WH13]. Market-Based [WH13]. Markets [BMV21, GSS16, BGMS11, DDGR11]. Massive [RCW22]. Matching [LST22, MHT19]. Mathematical [RRN21].

Matilda [KC13]. Matrix [LCS^{+}20]. Matter [WH13]. Maximization [MH13, WEM^{+}13]. means [LKZ^{+}15]. Measure [MFBK^{+}17]. Measurements [MPBH20]. measures [MCTC11]. Measuring [JJ19]. MEC [HDZ^{+}23]. MEC-enabled [HDZ^{+}23]. Mechanism [CWY^{+}23, HYYY22, LKZ^{+}15, LYC16, MHC^{+}22]. Media [AMZ20, AFD^{+}21, CRW19, LZLL21, LC19, MHC^{+}22, YYJ14, YZJ^{+}19, ZLC12]. Mediated [VSRU13, BFV12]. Medical [CLX^{+}23, EYS^{+}17, MFBK^{+}17, RRN21, XZBZ21, XZM^{+}22, ZML^{+}13, vdLSFEA23]. Medicare [MH13]. Medication [SXH^{+}24]. MediCoSpace [vdLSFEA23]. Meeting [RSM^{+}15]. Mental [BMR^{+}21]. Message [DMJ^{+}13]. Meta [RCFT18]. Meta-Model [RCFT18]. metaphor [Cha11]. Method [ALGA24, DZP^{+}24, FKBW15, NYS12, SCR23]. Methods [LFS18, MSD23, SB24]. metrics [ZLC12]. Microblog [PFD^{+}13]. Microblogging [BR15]. Miner [WWL^{+}22]. Minimizing [RSM^{+}15]. Minimum [BMRW15]. Mining [ARLEG17, ALGA24, BMFL15, CAL22, CSSD22, DLN^{+}21, EFPS^{+}22, GLC^{+}20, HBK^{+}20, KKB^{+}22, LSSD22, NFVY^{+}22, NYS12, NLLZ20, PWS^{+}15, PLW^{+}17, SXH^{+}24, WWL^{+}22, WWL^{+}24, YHLW15, ZZ23, ZML^{+}13, vdA12, FCKG10, LLK^{+}11]. Mismatch [LHT15]. Missing [HLG14]. Mitigation [YRUP14]. Mixed [RCFT18, SCR23]. Mixed-Effects [SCR23]. MOB [SCT^{+}13]. Mobile [LZW^{+}19, LZK21]. Modality [NJC^{+}22]. Model [FBP24, GH13, LCWH24, PCK21, RB23, RCFT18, VSRU13, WCL^{+}21, WZZ22, XLLX18, ZZA^{+}13, ZBX^{+}21, ZKMK21]. Modeling [CRAH10, DBC^{+}14, HMN^{+}16, JHC19, LTC^{+}12, LS17, NST21, PBT18, SCR23, WEM^{+}13, LLK^{+}11, PZH11, ALJH12]. Models [CRW19, KTF15, KM23, TWC16, FCKG10]. Moderating [OUAC21]. Modern [PPC^{+}20, PMH^{+}20]. Modular [PPG^{+}23]. Monitor [RCM^{+}22]. Monitoring [EV13, FKK^{+}23, LSM^{+}21]. monotonic [LL24]. MOOCs [HL21]. Mouth [ZGG13, ZLC12]. movement [LLKC11]. Much [SWB19]. Multi [CLX^{+}23, CWY^{+}23, HLYW22, JNDZ23, QGRT21, SKC20, SKC21, SG21]. Multi-attention [HLYW22]. Multi-dimensional [CWY^{+}23]. Multi-Disciplinary [SKC20, SKC21]. Multi-disease [QGRT21]. Multi-instance [SG21]. Multi-Label [CLX^{+}23]. Multi-Scale [JNDZ23]. Multichannel [LP16]. multilabel [HL11]. multimethod [Kan11]. Multimorbidity [SXH^{+}24, SJ23]. Multiple [DXC23, ZTD^{+}13, ZNJ19]. Mutual [ZJT^{+}23].

Name [Ema19]. Nation [Goo14]. Natures [JHC19]. Near [FKK^{+}23]. Near-Infrared [FKK^{+}23]. Need [ULST23]. Negotiation [HMN^{+}16]. Net [JNDZ23]. Netflix [GUH16]. Network [ALJH12, CXSW17, CFH22, CXG^{+}22, LKZ^{+}15, NSK^{+}21, NLLZ20, PPC^{+}20, TZLX21, WCL^{+}21, WR12, YDS^{+}13, ZTD^{+}13, WXR10]. Network-based [WCL^{+}21]. Networks [AGL23, BR15, CRW19, JFI^{+}23, LFS18, Orm13, PPC^{+}20, SMY20, ZGG13]. Neural [CXSW17]. News [AFD^{+}21, BR15, GHK^{+}20, GVP21, LTC^{+}12, ZZ23, Uhl11, BR15]. Next [LSS15]. Nodal [CRW19]. Noise [RCM^{+}22]. Non [LL24, MCCC22, MSU^{+}20]. Non-Compliance [MCCC22]. Non-monotonic [LL24]. Non-sensitive [MSU^{+}20]. nonroutine [DFP11]. Notes [BMFL15]. nothing [BKSZ10]. Novel [AHS^{+}23, LLF^{+}22]. Nursing [HYG19].

Object [CJH22]. Oblivious [PPC^{+}20].

Occupancy [LSM^{+}21]. occurrence [WWL^{+}24]. OFDM [LSM^{+}21]. OFDM-Based [LSM^{+}21]. off [COS^{+}21, KM23, WWL^{+}22]. Old [Tuz11]. Older [KC13]. One [LZW^{+}19, ULST23, WWL^{+}22, LCS^{+}20]. One-Handed [LZW^{+}19]. One-off [WWL^{+}22]. Online [BMR^{+}21, HCLN12, HBL14, KZ19, LSS20, LZX16, MSD17, NYS12, TZ20, WWL^{+}15, WZZS20, ZJT^{+}23, ZGG13, BFV12, LLK^{+}11]. Only [Bur16]. Ontology [LHT15, MHT19, XWC21]. Ontology-Based [LHT15]. Open [BDLC20, BNSW15, BSA14, KJRD16, WHEC22]. Open-Access [WHEC22]. Open-Source [BNSW15, BSA14]. Operations [KOH^{+}21, PHCS11]. Opportunities [KCSTM18, MWV^{+}18, vdA12]. Opportunity [CWY^{+}23, CGS12]. Optimal [LT16, PGPB19, RSM^{+}17, RSM^{+}21, SK12]. Optimization [LZX^{+}22a, RCW22]. Optimizing [BKMK23, HYG19, XWC21]. Options [KS13, DDGR11]. Order [SCT^{+}13, WWL^{+}24]. Order-preserving [WWL^{+}24]. Organizational [JADT15, RSM^{+}15, TZLX21, ALGA24]. Organizations [RSM^{+}21, KY11]. Organizing [WR12]. Oriented [YHLW15, ETHJ22, ZNJ19]. Orienteering [PPG^{+}23]. Originating [ALJH12]. Outages [BTZ17]. Outcomes [BMRW15, SXH^{+}24, ZTD^{+}13]. Outer [SB24]. Outlier [BDLC20, KKB^{+}22]. Outliers [BDD^{+}21]. Outsourced [MSU^{+}20]. Overcoming [AVBC23]. Overview [vdA12]. OWSP [WWL^{+}22]. OWSP-Miner [WWL^{+}22].

P2P [Sie23]. PANDA [MSU^{+}20]. Pandemic [KOH^{+}21, KJM23, KSS14, SJ23, ARV11]. Pandemics [GVP21, ZZYT21]. Papers [JRYY19]. Paradigm [EDBK21].

Paradigms [WZW22]. Part [CSSD22, LSSD22]. Participation [KJRD16]. Particle [LZX^{+}22a]. Partitioned [MSU^{+}20]. Partnerships [GHK^{+}20]. party [GHK^{+}20]. Passive [LSM^{+}21]. Paths [LL24]. Pathways [ZML^{+}13]. Patient [EFK^{+}20, SXH^{+}24, ZML^{+}13, vdLSFEA23, RADS12]. Patients [ABE^{+}21, HLG14, XZM^{+}22]. Pattern [CSSD22, DLN^{+}21, KKB^{+}22, LSSD22, LGW^{+}22, LZZ^{+}22, WWL^{+}22, WWL^{+}24]. Pattern-Driven [CSSD22, LSSD22]. patterns [AC10, LR11]. Paying [SWB19]. Payment [JFI^{+}23, Sre20]. Peer [Sie23]. Peer-to-Peer [Sie23]. Penny [LC19]. People [KC13]. Perceived [HBL14, Kak17]. Performance [DBC^{+}14, HSK17, LZX^{+}22b, LP16, ZZ23, AZ12]. Performance-Driven [LZX^{+}22b]. Peripheral [KJRD16]. permission [BKSZ10]. Person [Ema19, ZZX^{+}18]. Person-Job [ZZX^{+}18]. Personality [SZX^{+}23]. Personalization [LSS15]. Personalized [FN18, HBK^{+}20, PPG^{+}23]. Perspective [AY22, AVR21, EFK^{+}20, GSV15, HL21, JL15, MPBH20, SKC21, SKT15, WHEC22, HC11]. Phenomenon [SCT^{+}13]. Phone [LZK21]. Physical [GYZY22, LCF^{+}22]. PIE [AVR21]. Placement [DTL^{+}22]. Plans [MOR22]. Platform [MSD17, RCW22, SJKP17]. Platforms [Sie23]. Policies [RRN21, ZB21]. Policy [LZLL21, Sre20]. Portfolios [BMS17]. Position [SZX^{+}23]. Post [LKPZ23, PFD^{+}13, SXH^{+}24, FBP24, LSS20]. Post-Hip [SXH^{+}24]. Posting [WZZS20]. Postmarketing [YYJ14]. Power [BTZ17, HSK17]. Powered [HDZ^{+}23]. Practice [KF19]. Practices [JCMR15, RCM^{+}22]. Precision [LS17]. Predict [SXH^{+}24]. Predicting [AGL23, BSA14, WZZS20]. Prediction [CXG^{+}22, CSSD22, HLYW22, HLG14, HZZ12, LSSD22, MHC^{+}22, MSD23, TZLX21, TQ14, WCL^{+}21, XZM^{+}22, ZCL15, LLKC11]. Prediction-Driven [HZZ12]. Predictive [HBK^{+}20, QGRT21]. Predictors [MCCC22]. Presence [RR20, XLZ^{+}16]. Presented [JRYY19]. Preservation [CLYH19]. Preserving [CWY^{+}23, MO13, PGS22, WWL^{+}24]. Pressure [HLYW22]. Price [CSS17, XLLX18, LLKC11]. Pricing [HCLN12, BKSZ10]. Principles [ARLEG17, MMW11]. Prioritization [EFK^{+}20]. Privacy [AVR21, CWY^{+}23, EV13, EFPS^{+}22, EFK^{+}20, KLL19, MO13, PCK21, PGS22, RRN21, SKC21, SG21, ZB21, FCKG10]. Privacy-Enhanced [EV13]. Privacy-Preserving [CWY^{+}23, MO13, PGS22]. Pro [XZBZ21]. Proactive [ZELC22]. probabilistic [LLK^{+}11]. Problem [JBEM15]. Problems [DLN^{+}21, LHT15, SRS11]. Procedures [OUAC21]. Process [ALGA24, CCWL22, DFL^{+}23, EFPS^{+}22, GSV15, HBL14, KTF15, MWV^{+}18, OAHY21, PWS^{+}15, SK12, SJC15, SSY16, XLZ^{+}16, ZLY^{+}15, vdA12]. Processes [DM15, PWS^{+}15, WWL^{+}15, YHLW15]. Processing [POtH22, RCW22]. Product [JCMR15, Kak17, KZ19, QvdHT24, TLK20, ZGG13]. Product-service [QvdHT24]. productivity [HC11]. Products [NLLZ20]. Profile [LZK21]. Profiles [PLW^{+}17]. Program [CGS12, MH13]. Progress [BMFL15, ZB21]. Project [SJC15]. Projects [FKBW15, KJRD16, NM15]. Promise [ZB21]. Promoting [ALJH12]. propagation [ARV11]. Property [ACC^{+}21]. protection [FCKG10]. Protocol [PGS22]. prototype [MMW11]. prototypes [LLKC11]. Provenance [WRZ22]. PSO [LKZ^{+}15]. Psycholinguistics [DZP^{+}24]. Psycholinguistics-inspired [DZP^{+}24]. Public [LZLL21, MRKE23, DW11].

Publicly [YYJ14]. publishing [FCKG10]. Purchase [CSS17]. Push [LT16]. Putting [KF19].

Q&A [WWL^{+}15]. QoS [HMN^{+}16]. Qualities [HH19]. Quality [BMFL15, CLYH19, RCFT18, Kan11, LR11, PHCS11]. Quantile [LST22]. Quantile-Matching [LST22]. Quantitative [ETHJ22, FKBW15]. Query [TCS22]. Question [KF19]. Question-Answering [KF19].

R&D [SJC15]. Radiography [ZKMK21]. Random [ZZA^{+}13]. Randomized [RR20]. Rank [LCS^{+}20]. Rank-one [LCS^{+}20]. Rapid [GVP21]. Rate [GSS16]. Rating [HYG19, LTC^{+}12]. Ratings [AGL23, RR20]. Ratio [LT16]. Ratio-Optimal [LT16]. Ratios [LTC^{+}12]. Re [PPC^{+}20, PHL^{+}21]. Re- [PHL^{+}21]. Reactions [YRUP22]. Read [ZZ23]. Readmission [XZM^{+}22]. Real [GHK^{+}20, KS13, KTF15, KM23, LYC16]. Real-Time [KTF15, LYC16]. Realisation [OAHY21]. realities [KY11]. recall [BFV12]. Reciprocity [LSS20]. Recommend [HLCY22]. Recommendation [LFS18, ZZA^{+}13]. Recommendations [UTL20]. Recommender [GUH16, JJ19, AZ12]. Reconciliation [RRN21]. Reconstructing [ULST23]. Record [ZML^{+}13]. Recruitment [RSM^{+}21]. Red [HYG19]. Red-Handed [HYG19]. Redactable [MRKE23]. redesigning [MMW11]. Reflect [HCLN12]. Regional [LCWH24]. Regressional [NYS12]. Regularity [LZK21]. Regulatory [HYYY22, NM15]. Reinforcement [BKMK23]. Related [ZZYT21, ZCL15, DFP11, WCL^{+}21]. Relational [CXG^{+}22]. Relationship [LC19, NST21]. Relationships [BMRW15, Kak17]. Relay [STASH20]. ReliefF [NYS12]. ReliefF-Enhanced [NYS12]. Repetition [EYS^{+}17]. Reports [GVP21, LLKC11]. Repositories [WHEC22]. Representation [ZZX^{+}18]. representative [LLKC11]. Reputation [RR20, SNS^{+}20]. Requirement [RSM^{+}15]. Requirements [BMRW15, BNSW15, Bur16, CRD15, EFK^{+}20, FKBW15, JCMR15, JL15, JBEM15, KS15, NM15, TW17, SRS11]. Research [DFL^{+}23, EFPS^{+}22, GH13, HS23, LTL23, LFS18, LCC13, MOR22, SKC20, SB24, FCKG10]. researcher [HC11]. Reservation [CSS17]. Residential [KC13]. Resolution [ZBX^{+}21]. Resolving [DSZ17]. Resource [HDZ^{+}23, PLW^{+}17, XLZ^{+}16]. Resources [NHL22]. Response [VSRU13]. Responses [LZLL21]. Responsible [WRZ22]. results [Cha11]. Retail [KS13]. retention [PHCS11]. Return [ZCL15]. returns [DFP11]. Retweeting [ALJH12]. Revenue [SWB19]. Revenues [MSD17]. Review [CLYH19, NYS12, UL13, ZAZC18, LLK^{+}11]. Reviews [HCLN12, KZ19, TZ20, ZLC12]. revisited [Tuz11]. RFID [KS13, KTF15]. RFID-Based [KTF15]. Right [ZZX^{+}18]. Rights [CN23]. Risk [DDGR11, KSS14, PGPB19, PHL^{+}21, Sie23, YRUP14, HL11]. Roadmap [SKC20]. Roadside [CJH22]. Robot [KC13]. Robust [ABE^{+}21, PGPB19]. Role [BNSW15, DSZ17, GBK23, GSS16, KZ19, PCLL19, RSM^{+}17, TW17, ZLY^{+}15]. Role-Based [RSM^{+}17, ZLY^{+}15]. roles [DFP11]. Routing [LZX^{+}22b, LZX^{+}22a]. RT [BR15]. Rule [HBK^{+}20, MSD17]. Rule-Based [MSD17]. Rumors [AFD^{+}21].

Safe [LSM^{+}21]. Safety [YYJ14]. Samples [LSM^{+}21]. Satellite [KM23]. Scalable [OAHY21, PFD^{+}13]. Scale [FKBW15, JNDZ23, MPBH20, STBI22]. Scans [ABE^{+}21, JNDZ23]. Scheduling [XLZ^{+}16, LSS15]. Schemes [ETHJ22]. Science [GH13, LTL23, SB24, ZZYT21, Che11, NM12]. Scientific [DTL^{+}22]. Scoring [SNS^{+}20]. Search [GSS16, LCC17, WZW22, Cha11, PZH11, WXR10]. Searching [TQ14]. SEC [HL11]. Section [JRYY19, ZZYT21]. Sector [RCFT18, DW11]. Secure [MRKE23, XWLC19]. Security [EFK^{+}20, Goo14, LCF^{+}22, MSU^{+}20, OUAC21, RSM^{+}15, WHEC22, YRUP14, YRUP22, WXR10]. Segment [LZK21]. Segmented [SCR23]. Selection [BKMK23, CLYH19, MSD17, NSK^{+}21, RASHD22, SJC15]. Self [WWL^{+}22]. Self-adaptive [WWL^{+}22]. Seller [SJKP17]. Semantic [KZ19]. Semi [AHS^{+}23]. Semi-automated [AHS^{+}23]. sense [SKT11]. Sensemaking [CRD15]. Sensing [PPG^{+}23]. Sensitive [MSU^{+}20, PFD^{+}13]. Sensors [BTZ17]. Sentiment [CXG^{+}22, DSZ17, DXC23, ZAZC18, Uhl11]. Septic [HLG14]. Sequences [MFBK^{+}17]. Sequential [HBK^{+}20]. Serendipity [FN18]. Series [MSD23, WEM^{+}13, WWL^{+}24, AC10]. Server [HSK17]. Service [HMN^{+}16, LT16, LSS15, LP16, PPC^{+}20, ZNJ19, QvdHT24, PHCS11]. Service-oriented [ZNJ19]. Services [DJS18, EYS^{+}17, LSS15]. Set [CLYH19]. Shalls [CRD15]. Shanghai [WCL^{+}21]. Sharing [JFI^{+}23, LSS20, WWL^{+}15]. Shock [HLG14, XLLX18]. Shots [LST22]. Side [MSD17]. Sight [LZW^{+}19]. Sight-Free [LZW^{+}19]. Signals [EDBK21]. Similarities [TQ14]. Similarity [MFBK^{+}17]. Simon [SB24]. Simulated [MHT19, NFVY^{+}22]. Simulating [FBP24]. Situated [JCMR15]. Situational [CN23, HBK^{+}20]. Six [MMW11]. Small [PGPB19]. Smart [CCWL22, CTED22, CJH22, KCSTM18, LGW^{+}22, LZX^{+}22b, LZX^{+}22a, LCF^{+}22, NHL22, RCW22, YLA13]. SOA [CNJ13]. Social [AFD^{+}21, BTZ17, CRW19, GYZY22, GBK23, LFS18, LZLL21, LC19, MHC^{+}22, PPG^{+}23, Sie23, WR12, XLLX18, YYJ14, YZJ^{+}19, ZTD^{+}13, ZZA^{+}13, KY11, PZH11, RW11, ZLC12]. social-broadcasting-based [RW11]. societal [Che11]. Socioeconomic [PCLL19]. Software [BNSW15, BSA14, JHC19, Kak17, CRAH10]. Solutions [PBT18]. Solving [DLN^{+}21]. something [BKSZ10]. Source [BNSW15, BSA14, Bur16, KJRD16]. Sources [Bur16, DXC23]. Space [NST21, TWC16]. Spaces [vdLSFEA23]. spam [LLK^{+}11]. Spanning [JCMR15]. Sparse [WZZ22]. Spatial [LSM^{+}21]. Spatio [ZBX^{+}21]. Spatio-Temporal [ZBX^{+}21]. Special [BJJK17, CTED22, CSSD22, DFJ20, JRYY19, LTL23, LKPZ23, LSSD22, ZZYT21]. specific [MCTC11]. Spectroscopy [FKK^{+}23]. Sponsored [GSS16, LCC17]. Spyware [PMH^{+}20]. Stability [BSA14]. Stacked [WZZ22]. Stage [LZX^{+}22a]. Stakeholder [ZCL15]. Stakeholders [Bur16]. State [ZAZC18]. State-of-the-Art [ZAZC18]. States [SYS^{+}13]. Status [PCLL19]. Stock [LLKC11, TQ14, ZCL15]. Stocks [LC19]. Storage [MRKE23, RCW22, DDGR11]. Strategies [LZX16, BKSZ10]. Strategy [DTL^{+}22, SJKP17]. Streams [PFD^{+}13, MAKH^{+}11, RADS12]. Strong [WWL^{+}22]. Structural [BNSW15, SKT15]. Study [BMFL15, BNSW15, ETHJ22, KTF15, NM15, RCFT18, Sre20, WCL^{+}21, Kan11]. Subgraphs [CAL22]. subject [PZH11]. subject-centered [PZH11]. Substance [BKMK23]. Supervised [TZ20]. Supply [KTF15, MSD17]. Supply-Side [MSD17]. Support [ARLEG17, LZW^{+}19, LLF^{+}22, MO13, vdLSFEA23, ARV11, KY11]. Supported [NMS12]. Supporting [CXG^{+}22]. Surveillance [YYJ14]. Survey [LFS18, MCCC22, MMW11]. Survey-based [MCCC22]. Survivability [NHL22]. sustainability [AC10]. Swarm [LZX^{+}22a].

Swine [KSS14]. Symptom [GVP21]. SymptomID [GVP21]. Synchronous [SK12]. Synthesized [SMY20]. System [AHS^{+}23, ABE^{+}21, CNJ13, DJS18, GLC^{+}20, GUH16, HYG19, KS13, LGW^{+}22, LP16, MO13, PPG^{+}23, RR20, RCM^{+}22, WCX21, YZJ^{+}19, ZML^{+}13, RW11]. Systematic [UL13]. Systemic [GSS16]. Systems [ARLEG17, AHS^{+}23, BSA14, CTED22, CGS12, CMM^{+}22, DFL^{+}23, GYZY22, HS23, HDZ^{+}23, JRYY19, JJ19, JL15, JBEM15, KCSTM18, KF19, LT16, LCF^{+}22, LLF^{+}22, NM15, OAHY21, PBT18, RSM^{+}17, RB23, RCFT18, SKC21, SSY16, VSRU13, WRZ22, ZZA^{+}13, AZ12, HC11, KY11, MMW11, NM12, NB11]. systems-findings [MMW11].

Tackling [KM23]. Tagging [ZZA^{+}13, PZH11]. Tags [WR12]. Tale [PHL^{+}21]. Talent [ZZX^{+}18]. Target [CXG^{+}22]. Targeted [DXC23]. Targeting [LZX16]. Task [SKT15]. Tasks [HDZ^{+}23, DFP11]. Taxi [HLCY22]. Taxonomies [BDLC20]. Taxonomy [Bur16]. Teams [SKT15, BFV12, MM12]. Technical [AY22]. Technique [LHT15, LYC16, FCKG10]. Technologies [OUAC21]. Technology [HYYY22, JRYY19, KS13, UL13, YRUP22, ZTD^{+}13, BFV12, MCTC11]. technology-mediated [BFV12]. tele [SRS11]. tele-advisory [SRS11]. telesales [SRS11]. Temporal [CRW19, ZBX^{+}21]. term [ZZ23]. Terms [TZ20]. Terraform [RCM^{+}22]. Test [EDBK21, SYS^{+}13]. testing [DW11]. Text [ARLEG17, BMFL15, LZW^{+}19, LLK^{+}11, NYS12, NJC^{+}22, ZZ23, ZLC12, ZELC22, HL11]. Text-Based [ZELC22, ZLC12]. Textual [Sie23]. Theft [ACC^{+}21, DZP^{+}24]. Their [CSS17, LR11, MM12]. Them [HYG19]. Theoretical [LFS18, SB24]. Theorization [KJM23]. Theory [PPC^{+}20]. Things [ETHJ22, MPBH20]. Third [GHK^{+}20].

Third-party [GHK^{+}20]. Thousand [LC19]. Threat [ZELC22]. Threats [EFPS^{+}22]. Three [WZW22]. Threefold [HH19]. Thrive [PHL^{+}21]. Throughput [WEM^{+}13]. ThumbStroke [LZW^{+}19]. Time [AMZ20, GSS16, KTF15, LYC16, MSD23, WEM^{+}13, WWL^{+}24, AC10]. Time-based [AMZ20]. time-series [AC10]. Times [LSS15]. Timing [LZX^{+}22a]. Timing-Driven [LZX^{+}22a]. TMIS [BJJK17, Che10, DFJ20]. Token [FBP24, PZW24]. Token-Based [PZW24]. too [SWB19]. Tool [KSS14]. Toolkit [KOH^{+}21]. Tools [ZJT^{+}23]. Topics [PFD^{+}13]. Touchscreen [LZW^{+}19]. Toulmin [RB23]. Tracing [AVBC23]. Trade [COS^{+}21, KM23]. Trade-off [COS^{+}21, KM23]. Traffic [NSK^{+}21]. Trailblazing [SKC20]. Trajectories [SXH^{+}24]. Trajectory [BDLC20, BDD^{+}21, XZM^{+}22]. Trajectory-Based [XZM^{+}22]. Transaction [TSCT18]. Transfer [KF19, NJC^{+}22]. transitions [RADS12]. Translation [TLK20]. Transparency [NMS12]. travel [SRS11]. Treatment [BKMK23]. Tree [LZZ^{+}22]. Tree-based [LZZ^{+}22]. Trend [GLC^{+}20]. Trending [PFD^{+}13]. Trends [AMZ20, HYYY22]. TRG [CXG^{+}22]. TRG-DAtt [CXG^{+}22]. tricks [Tuz11]. Trust [CSS17, ETHJ22, HBL14, LCC17, MCTC11, Orm13, RB23]. Trustable [CMM^{+}22]. Truthful [SWB19]. Turnover [TZLX21]. Tweeters [ALJH12]. Twins [LCF^{+}22]. Twitter [ALJH12, KSS14, ZAZC18]. Two [HZZ12, KJM23, LZX^{+}22a]. Two-Stage [LZX^{+}22a]. Typing [DMJ^{+}13].

U.S [HL21, ZBX^{+}21]. U.S. [SJ23, Uhl11]. Uncovering [DW11, KY11]. Understand [AY22]. Understanding [BRH^{+}15, BMS17, BMR^{+}21, LL24, WWL^{+}15]. Unfair [RR20]. Unique [CGS12]. Unit [CJH22].

Unit-based [CJH22]. Universal [SSY16]. Universities [Goo14]. University [KOH^{+}21]. Unknown [CJH22]. Untrustworthy [GHK^{+}20]. Urban [LCWH24]. Use [AVBC23, BKMK23, Kak17, KSS14, BFV12, DFP11, MM12]. User [Kak17, LZK21, PGS22, RSM^{+}15, WZZS20, WHEC22]. User-empowered [PGS22]. Using [ACC^{+}21, AGL23, ABE^{+}21, BMS17, BSPP18, BKMK23, BTZ17, BSA14, CMM^{+}22, DZP^{+}24, DMJ^{+}13, GHK^{+}20, GVP21, JFI^{+}23, LKZ^{+}15, LSM^{+}21, LZLL21, LTC^{+}12, MOR22, MSD23, NYS12, PCK21, RB23, Sie23, WR12, YYJ14, ZZYT21, Cha11, LLKC11, MHC^{+}22, vdLSFEA23]. Utility [BTZ17, GLC^{+}20, GH13, LGW^{+}22, NFVY^{+}22]. Utility-Driven [GLC^{+}20]. Utilization [GBK23].

Valuation [WH13]. Value [CSS17, FKBW15, GUH16, JJ19, DDGR11]. Values [Kak17]. Variable [LSS15]. Variety [SMY20]. Vast [KLL19]. Vehicle [PGS22]. Vehicles [CJH22]. Vehicular [CFH22]. Vendor [HCLN12, SJKP17]. Veracity [SMY20]. Verifiable [PGS22]. via [HSK17, RASHD22, TLK20, TWC16, WEM^{+}13, ZZ23, ZML^{+}13]. Video [EV13]. View [ZLY^{+}15]. Violence [AVBC23]. Virtual [LZW^{+}19, LAYR13, BKSZ10, MM12, SKT11]. vision [NB11]. Visits [SJ23]. Visual [BMS17, HH19, vdLSFEA23, BFV12]. Visualization [BSPP18]. Visualizing [Cha11]. vs [NMS12, SK12]. vulnerability [WXR10].

wage [DFP11]. Walk [ZZA^{+}13]. Warning [KSS14]. Waves [KJM23]. Way [LST22]. Weak [WWL^{+}22]. Weak-gap [WWL^{+}22]. Weakly [TZ20]. Wearables [SCR23]. Weather [CJH22]. Web [Cha11, DJS18, DBC^{+}14, Ema19, HMN^{+}16, WR12, ZELC22, ZCL15]. Web-Based [DBC^{+}14]. Website [HH19]. Websites [GHK^{+}20]. Weighted [LCS^{+}20]. Welcome [Che10]. Wellbeing [KC13, YLA13]. while [RSM^{+}15]. Who [ALJH12, LR11, LSS15]. Wi [JFI^{+}23]. Wi-Fi [JFI^{+}23]. WiFi [LSM^{+}21]. wiki [Kan11]. Wikipedia [LR11]. wikis [AC10]. Will [PHL^{+}21]. Window [GSS16]. Wirelessly [HDZ^{+}23]. within [BFV12]. WITS [BJJK17, DFJ20]. Word [ACC^{+}21, LHT15, ZGG13, ZLC12]. Word-of-Mouth [ZGG13, ZLC12]. WordNet [TZ20]. WordNet-Guided [TZ20]. Work [SZX^{+}23]. Workflow [TCS22]. Workflows [DTL^{+}22]. Workforce [Goo14]. Workshop [JRYY19]. workspace [MM12]. World [DAHND18]. Worth [LC19]. Write [XZBZ21].

X [LZX^{+}22b, LZX^{+}22a]. X-Architecture [LZX^{+}22b]. X-Routing [LZX^{+}22a].

References

Avetisian:2021:CSR

[ABE^{+}21] Manvel Avetisian, Ilya Burenko, Konstantin Egorov, Vladimir Kokh, Aleksandr Nesterov, Aleksandr Nikolaev, Alexander Ponomarchuk, Elena Sokolova, Alex Tuzhilin, and Dmitry Umerenkov. CoR-SAI: a system for robust interpretation of CT scans of COVID-19 patients using deep learning. ACM Transactions on Management Information Systems (TMIS), 12(4):28:1–28:16, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                                    //dl.acm.org/doi/10.1145/
                                                    3467471.

Arazy:2010:SCW

[AC10] Ofer Arazy and Arie Croitoru.

The sustainability of corporate wikis: a time-series analysis of activity patterns. ACM Transactions on Management Information Systems (TMIS), 1(1): 6:1–6:??, December 2010. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Abdibayev:2021:UWE

[ACC^{+}21] Almas Abdibayev, Dongkai Chen, Haipeng Chen, Deepti Poluru, and V. S. Subrahmanian. Using word embeddings to deter intellectual property theft through automated generation of fake documents. ACM Transactions on Management Information Systems (TMIS), 12(2):13:1–13:22, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3418289.

Alkhodair:2021:DHE

[AFD^{+}21] Sarah A. Alkhodair, Benjamin C. M. Fung, Steven H. H. Ding, William K. Cheung, and Shih-Chia Huang. Detecting highengaging breaking news rumors in social media. ACM Transactions on Management Information Systems (TMIS), 12(1): 8:1–8:16, March 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3416703.

Ang:2023:PER

[AGL23] Gary Ang, Zhiling Guo, and Ee-Peng Lim. On predicting ESG ratings using dynamic company networks. ACM Transactions on Management Information Systems (TMIS), 14(3):27:1–27:??, September 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3607874.

Ameri:2023:DNI

[AHS^{+}23] Kimia Ameri, Michael Hempel, Hamid Sharif, Juan Lopez, and Kalyan Perumalla. Design of a novel information system for semi-automated management of cybersecurity in industrial control systems. ACM Transactions on Management Information Systems (TMIS), 14(1): 4:1–4:??, March 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3546580.

Aouachria:2024:PMM

[ALGA24] Moufida Aouachria, Abderrahmane Leshob, Abdessamed Réda Ghomari, and Mustapha Aouache. A process mining method for inter-organizational business process integration. ACM Transactions on Management Information Systems (TMIS), 15(1):2:1–2:??, March 2024. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638062.

Achananuparp:2012:WRT

[ALJH12] Palakorn Achananuparp, Ee-Peng Lim, Jing Jiang, and Tuan-Anh Hoang. Who is retweeting the tweeters? Modeling, originating, and promoting behaviors in the Twitter network. ACM Transactions on Management Information Systems (TMIS), 3(3):13:1–13:??, October 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Ahmed:2022:HEA

[ALS22] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Heterogeneous energyaware load balancing for Industry 4.0 and IoT environments. ACM Transactions on Management Information Systems (TMIS), 13(4):46:1–46:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3543859.

Alagheband:2020:TBG

[AMZ20] Mahdi R. Alagheband, Atefeh Mashatan, and Morteza Zihayat. Time-based gap analysis of cybersecurity trends in academic and digital media. ACM Transactions on Management Information Systems (TMIS), 11(4):20:1–20:20, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3389684.

Al-Ramahi:2017:DDP

[ARLEG17] Mohammad A. Al-Ramahi, Jun Liu, and Omar F. El-Gayar. Discovering design principles for health behavioral change support systems: a text mining approach. ACM Transactions on Management Information Systems (TMIS), 8(2–3): 5:1–5:??, August 2017. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Arora:2011:DSC

[ARV11] Hina Arora, T. S. Raghu, and Ajay Vinze. Decision support for containing pandemic propagation. ACM Transactions on Management Information Systems (TMIS), 2(4):23:1–23:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Akello:2023:BUC

[AVBC23] Patricia Akello, Naga Vemprala, Nicole Lang Beebe, and Kim-Kwang Raymond Choo. Blockchain use case in ballistics and crime gun tracing and intelligence: Toward overcoming gun violence. ACM Transactions on Management Information Systems (TMIS), 14(1): 7:1–7:??, March 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3571290.

Akanfe:2021:DIF

[AVR21] Oluwafemi Akanfe, Rohit Valecha, and H. Raghav Rao. Design of an inclusive financial privacy index (INF-PIE): a financial privacy and digital financial inclusion perspective. ACM Transactions on Management Information Systems (TMIS), 12(1):7:1–7:21, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3403949.

Adomavicius:2022:IBE

[AY22] Gediminas Adomavicius and Mochen Yang. Integrating behavioral, economic, and technical insights to understand and address algorithmic bias: a human-centric perspective. ACM Transactions on Management Information Systems (TMIS), 13(3):34:1–34:27, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3519420.

Adomavicius:2012:IDC

[AZ12] Gediminas Adomavicius and Jingjing Zhang. Impact of data characteristics on recommender systems performance. ACM Transactions on Management Information Systems (TMIS), 3 (1):3:1–3:??, April 2012. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Belhadi:2021:MLI

[BDD^{+}21] Asma Belhadi, Youcef Djenouri, Djamel Djenouri, Tomasz Michalak, and Jerry Chun-Wei Lin. Machine learning for identifying group trajectory outliers. ACM Transactions on Management Information Systems (TMIS), 12(2):12:1–12:25, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3430195.

Belhadi:2020:TOD

[BDLC20] Asma Belhadi, Youcef Djenouri, Jerry Chun-Wei Lin, and Alberto Cano. Trajectory outlier detection: Algorithms, taxonomies, evaluation, and open challenges. ACM Transactions on Management Information Systems (TMIS), 11(3): 16:1–16:29, August 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3399631.

Bartenschlager:2023:CML

[BEK^{+}23] Christina C. Bartenschlager, Stefanie S. Ebel, Sebastian Kling, Janne Vehreschild, Lutz T. Zabel, Christoph D. Spinner, Andreas Schuler, Axel R. Heller, Stefan Borgmann, Reinhard Hoffmann, Siegbert Rieg, Helmut Messmann, Martin Hower, Jens O. Brunner, Frank Hanses, and Christoph Römmele. COVIDAL: a machine learning classifier for digital COVID-19 diagnosis in German hospitals. ACM Transactions on Management Information Systems (TMIS), 14(2): 14:1–14:??, June 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3567431.

Basoglu:2012:ERW

[BFV12] K. Asli Basoglu, Mark A.

Fuller, and Joseph S. Valacich. Enhancement of recall within technology-mediated teams through the use of online visual artifacts. ACM Transactions on Management Information Systems (TMIS), 3(1):2:1– 2:??, April 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bhattacharjee:2011:DGM

[BGMS11] Sudip Bhattacharjee, Ram D.

Gopal, James R. Marsden, and Ramesh Sankaranarayanan. Digital goods and markets: Emerging issues and challenges. ACM Transactions on Management Information Systems (TMIS), 2(2):8:1–8:??, June 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bhattacharjee:2017:IWS

[BJJK17] Sudip Bhattacharjee, Varghese Jacob, Zhengrui (Jeffrey) Jiang, and Subodha Kumar. Introduction to WITS 2015 special issue in TMIS. ACM Transactions on Management Information Systems (TMIS), 8(2–3):4:1–4:??, August 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Baucum:2023:OSU

[BKMK23] Matt Baucum, Anahita Khojandi, Carole Myers, and Larry Kessler. Optimizing substance use treatment selection using reinforcement learning. ACM Transactions on Management Information Systems (TMIS), 14(2):13:1–13:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3563778.

Ba:2010:WGS

[BKSZ10] Sulin Ba, Dan Ke, Jan Stallaert, and Zhongju Zhang. Why give away something for nothing? Investigating virtual goods pricing and permission strategies. ACM Transactions on Management Information Systems (TMIS), 1(1):4:1– 4:??, December 2010. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Berndt:2015:CSD

[BMFL15] Donald J. Berndt, James A. McCart, Dezon K. Finch, and Stephen L. Luther. A case study of data quality in text mining clinical progress notes. ACM Transactions on Management Information Systems (TMIS), 6(1):1:1–1:??, February 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Biester:2021:UIC

[BMR^{+}21] Laura Biester, Katie Matton, Janarthanan Rajendran, Emily Mower Provost, and Rada Mihalcea. Understanding the impact of COVID-19 on online mental health forums. ACM Transactions on Management Information Systems (TMIS), 12(4):31:1–31:28, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458770.

Bai:2015:RAM

[BMRW15] Xue Bai, James R. Marsden, William T. Ross, Jr., and Gang Wang. Relationships among minimum requirements, Facebook likes, and Groupon deal outcomes. ACM Transactions on Management Information Systems (TMIS), 6(3):9:1–9:??, October 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Basole:2017:UAP

[BMS17] Rahul C. Basole, Timothy Major, and Arjun Srinivasan. Understanding alliance portfolios using visual analytics. ACM Transactions on Management Information Systems (TMIS), 8(4):12:1–12:??, September 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Benedetto:2021:EIC

[BMV21] Francesco Benedetto, Loretta Mastroeni, and Pierluigi Vellucci. Extraction of information content exchange in financial markets by an entropy analysis. ACM Transactions on Management Information Systems (TMIS), 12(1):9:1–9:16, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3419372.

Bhowmik:2015:RSH

[BNSW15] Tanmay Bhowmik, Nan Niu, Prachi Singhania, and Wentao Wang. On the role of structural holes in requirements identification: an exploratory study on open-source software development. ACM Transactions on Management Information Systems (TMIS), 6(3):10:1–10:??, October 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bhattacharya:2015:RNA

[BR15] Devipsita Bhattacharya and Sudha Ram. RT @News: an analysis of news agency ego networks in a microblogging environment. ACM Transactions on Management Information Systems (TMIS), 6(3): 11:1–11:??, October 2015. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Basole:2015:UBE

[BRH^{+}15] Rahul C. Basole, Martha G.

Russell, Jukka Huhtamäki, Neil Rubens, Kaisa Still, and Hyunwoo Park. Understanding business ecosystem dynamics: a data-driven approach. ACM Transactions on Management Information Systems (TMIS), 6 (2):6:1–6:??, July 2015. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bouktif:2014:PSO

[BSA14] Salah Bouktif, Houari Sahraoui, and Faheem Ahmed. Predicting stability of open-source software systems using combination of Bayesian classifiers. ACM Transactions on Management Information Systems (TMIS), 5(1):3:1–3:??, April 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Basole:2018:EDE

[BSPP18] Rahul C. Basole, Arjun Srinivasan, Hyunwoo Park, and Shiv Patel. ecoxight: Discovery, exploration, and analysis of business ecosystems using interactive visualization. ACM Transactions on Management Information Systems (TMIS), 9(2):6:1–6:??, September 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Bauman:2017:USS

[BTZ17] Konstantin Bauman, Alexander Tuzhilin, and Ryan Zaczynski. Using social sensors for detecting emergency events: a case of power outages in the electrical utility industry. ACM Transactions on Management Information Systems (TMIS), 8(2–3):7:1–7:??, August 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Burnay:2016:SOS

[Bur16] Corentin Burnay. Are stakeholders the only source of information for requirements engineers? Toward a taxonomy of elicitation information sources. ACM Transactions on Management Information Systems (TMIS), 7(3):8:1–8:??, October 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chowdhury:2022:NAM

[CAL22] Mohammad Ehsan Shahmi Chowdhury, Chowdhury Farhan Ahmed, and Carson K. Leung. A new approach for mining correlated frequent subgraphs. ACM Transactions on Management Information Systems (TMIS), 13(1):9:1–9:28, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473042.

Chen:2022:CIM

[CCWL22] Rongli Chen, Xiaozhong Chen, Lei Wang, and Jianxin Li. The core industry manufacturing process of electronics assembly based on smart manufacturing. ACM Transactions on Management Information Systems (TMIS), 13(4):40:1–40:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529098.

Chen:2022:ABV

[CFH22] Mu-Yen Chen, Min-Hsuan Fan, and Li-Xiang Huang. AIbased vehicular network toward 6G and IoT: Deep learning approaches. ACM Transactions on Management Information Systems (TMIS), 13(1): 6:1–6:12, March 2022. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3466691.

Chiang:2012:BIA

[CGS12] Roger H. L. Chiang, Paulo Goes, and Edward A. Stohr. Business intelligence and analytics education, and program development: a unique opportunity for the information systems discipline. ACM Transactions on Management Information Systems (TMIS), 3(3): 12:1–12:??, October 2012. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chau:2011:VWS

[Cha11] Michael Chau. Visualizing Web search results using glyphs: Design and evaluation of a flower metaphor. ACM Transactions on Management Information Systems (TMIS), 2(1):2:1–2:??, March 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chen:2010:EWF

[Che10]  Hsinchun Chen.  Editorial:

Welcome to the first issue of ACM TMIS. ACM Transactions on Management Information Systems (TMIS), 1(1):1:1– 1:??, December 2010. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chen:2011:EDS

[Che11] Hsinchun Chen. Editorial: Design science, grand challenges, and societal impacts. ACM Transactions on Management Information Systems (TMIS), 2 (1):1:1–1:??, March 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chen:2022:RUB

[CJH22] Yu-Chia Chen, Sin-Ye Jhong, and Chih-Hsien Hsia. Roadside unit-based unknown object detection in adverse weather conditions for smart Internet of Vehicles. ACM Transactions on Management Information Systems (TMIS), 13(4):47:1–47:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3554923.

Chai:2023:MLC

[CLX^{+}23] Yidong Chai, Hongyan Liu, Jie Xu, Sagar Samtani, Yuanchun Jiang, and Haoxin Liu. A multi-label classification with an adversarial-based denoising autoencoder for medical image annotation. ACM Transactions on Management Information Systems (TMIS), 14(2):19:1– 19:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561653.

Chen:2019:ESC

[CLYH19] Jiawei Chen, Hongyan Liu, Yinghui (Catherine) Yang, and Jun He. Effective selection of a compact and high-quality review set with information preservation. ACM Transactions on Management Information Systems (TMIS), 10 (4):15:1–15:??, December 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3369395.

Corradini:2022:ETA

[CMM^{+}22] Flavio Corradini, Alessandro Marcelletti, Andrea Morichetta, Andrea Polini, Barbara Re, and Francesco Tiezzi. Engineering trustable and auditable choreography-based systems using blockchain. ACM Transactions on Management Information Systems (TMIS), 13(3): 31:1–31:53, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3505225.

Chua:2023:SFD

[CN23] Cecil Eng Huang Chua and Fred Niederman. Situational factor determinants of the allocation of decision rights to edge computers. ACM Transactions on Management Information Systems (TMIS), 14(3): 24:1–24:??, 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3582081.

Choi:2013:ISI

[CNJ13] Jae Choi, Derek L. Nazareth, and Hemant K. Jain. The impact of SOA implementation on IT-business alignment: a system dynamics approach. ACM Transactions on Management Information Systems (TMIS), 4 (1):3:1–3:??, April 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chakraborty:2021:MLA

[COS^{+}21] Saurav Chakraborty, Agnieszka Onuchowska, Sagar Samtani, Wolfgang Jank, and Brandon Wolfram. Machine learning for automated industrial IoT attack detection: an efficiency-complexity trade-off. ACM Transactions on Management Information Systems (TMIS), 12(4):37:1–37:28, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460822.

Cao:2010:MDA

[CRAH10] Lan Cao, Balasubramaniam Ramesh, and Tarek Abdel-Hamid. Modeling dynamics in agile software development.

ACM Transactions on Management Information Systems (TMIS), 1(1):5:1–5:??, December 2010. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chakraborty:2015:GSF

[CRD15] Suranjan Chakraborty, Christoph Rosenkranz, and Josh Dehlinger. Getting to the shalls: Facilitating sensemaking in requirements engineering. ACM Transactions on Management Information Systems (TMIS), 5(3):14:1–14:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Chung:2019:IMD

[CRW19] Wingyan Chung, Bingbing Rao, and Liqiang Wang. Interaction models for detecting nodal activities in temporal social media networks. ACM Transactions on Management Information Systems (TMIS), 10(4):14:1–14:??, December 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3365537.

Cazier:2017:VCT

[CSS17] Joseph Cazier, Benjamin Shao, and Robert St. Louis. Value congruence, trust, and their effects on purchase intention and reservation price. ACM Transactions on Management Information Systems (TMIS), 8(4): 13:1–13:??, September 2017.

CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Cwei:2022:ISI

[CSSD22] Lin Jerry Cwei, Nachiketa Sahoo, Gautam Srivastava, and Weiping Ding. Introduction to the special issue on patterndriven mining, analytics, and prediction for decision making, Part 1. ACM Transactions on Management Information Systems (TMIS), 13(1):1:1–1:3, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486960.

Chen:2022:ISI

[CTED22] Mu-Yen Chen, Bhavani Thuraisingham, Erol Egrioglu, and Jose De Jesus Rubio. Introduction to the special issue on smart systems for Industry 4.0 and IoT. ACM Transactions on Management Information Systems (TMIS), 13(4):35:1–35:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3583985.

Chen:2023:ODD

[CWY^{+}23] Xue Chen, Cheng Wang, Qing Yang, Teng Hu, and Changjun Jiang. The opportunity in difficulty: a dynamic privacy budget allocation mechanism for privacy-preserving multidimensional data collection. ACM Transactions on Management Information Systems (TMIS), 14(1):8:1–8:??, March 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569944.

Chen:2022:TDT

[CXG^{+}22] Fan Chen, Jiaoxiong Xia, Honghao Gao, Huahu Xu, and Wei Wei. TRG-DAtt: The target relational graph and double attention network based sentiment analysis and prediction for supporting decision making. ACM Transactions on Management Information Systems (TMIS), 13(1):3:1–3:25, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462442.

Chen:2017:DLN

[CXSW17] Hao Chen, Keli Xiao, Jinwen Sun, and Song Wu. A doublelayer neural network framework for high-frequency forecasting. ACM Transactions on Management Information Systems (TMIS), 7(4):11:1–11:??, January 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

De-Arteaga:2018:MLD

[DAHND18] Maria De-Arteaga, William Herlands, Daniel B. Neill, and Artur Dubrawski. Machine learning for the developing world. ACM Transactions on Management Information Systems (TMIS), 9(2):9:1–9:??, September 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Deodhar:2014:IWB

[DBC^{+}14] Suruchi Deodhar, Keith R. Bisset, Jiangzhuo Chen, Yifei Ma, and Madhav V. Marathe. An interactive, Web-based high performance modeling environment for computational epidemiology. ACM Transactions on Management Information Systems (TMIS), 5(2):7:1– 7:??, July 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Du:2011:RHS

[DDGR11] Anna Ye Du, Sanjukta Das, Ram D. Gopal, and R. Ramesh. Risk hedging in storage grid markets: Do options add value to forwards? ACM Transactions on Management Information Systems (TMIS), 2(2): 10:1–10:??, June 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Dutta:2020:IWS

[DFJ20] Kaushik Dutta, Xiao Fang, and Zhengrui (Jeffrey) Jiang. Introduction to WITS 2018 special issue in TMIS. ACM Transactions on Management Information Systems (TMIS), 11(3): 9:1–9:2, August 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3404392.

Dumas:2023:AAB

[DFL^{+}23] Marlon Dumas, Fabiana Fournier, Lior Limonad, Andrea Marrella, Marco Montali, Jana-Rebecca Rehse, Rafael Accorsi, Diego Calvanese, Giuseppe De Giacomo, Dirk Fahland, Avigdor Gal, Marcello La Rosa, Hagen Völzer, and Ingo Weber. AI-augmented business process management systems: a research manifesto. ACM Transactions on Management Information Systems (TMIS), 14(1): 11:1–11:??, March 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3576047.

Dey:2011:CUW

[DFP11] Debabrata Dey, Ming Fan, and Gang Peng. Computer use and wage returns: The complementary roles of IT-related human capital and nonroutine tasks. ACM Transactions on Management Information Systems (TMIS), 2(1):6:1–6:??, March 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Davis:2010:IFF

[DGM^{+}10] Gordon B. Davis, Paul Gray, Stuart Madnick, Jay F. Nunamaker, Ralph Sprague, and Andrew Whinston. Ideas for the future of the IS field. ACM Transactions on Management Information Systems (TMIS), 1(1):2:1–2:??, December 2010.

CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Delano:2018:SDT

[DJS18] John D. Delano, Hemant K. Jain, and Atish P. Sinha. System design through the exploration of contemporary Web services. ACM Transactions on Management Information Systems (TMIS), 9(3):13:1– 13:??, November 2018. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3273932.

Djenouri:2021:EDS

[DLN^{+}21] Youcef Djenouri, Jerry Chun-Wei Lin, Kjetil Nørvåg, Heri Ramampiaro, and Philip S. Yu. Exploring decomposition for solving pattern mining problems. ACM Transactions on Management Information Systems (TMIS), 12(2):15:1–15:36, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3439771.

Ciccio:2015:DDC

[DM15] Claudio Di Ciccio and Massimo Mecella. On the discovery of declarative control flows for artful processes. ACM Transactions on Management Information Systems (TMIS), 5(4): 24:1–24:??, January 2015. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Derrick:2013:DDC

[DMJ^{+}13] Douglas C. Derrick, Thomas O.

Meservy, Jeffrey L. Jenkins, Judee K. Burgoon, and Jay F. Nunamaker, Jr. Detecting deceptive chat-based communication using typing behavior and message cues. ACM Transactions on Management Information Systems (TMIS), 4(2):9:1– 9:??, August 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Deng:2017:RAS

[DSZ17] Shuyuan Deng, Atish P. Sinha, and Huimin Zhao. Resolving ambiguity in sentiment classification: The role of dependency features. ACM Transactions on Management Information Systems (TMIS), 8(2–3):4:1–4:??, August 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Du:2022:SWI

[DTL^{+}22] Xin Du, Songtao Tang, Zhihui Lu, Keke Gai, Jie Wu, and Patrick C. K. Hung. Scientific workflows in IoT environments: a data placement strategy based on heterogeneous edge-cloud computing. ACM Transactions on Management Information Systems (TMIS), 13(4):42:1–42:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531327.

Dawson:2011:UTA

[DW11] Gregory S. Dawson and Richard T. Watson. Uncovering and testing archetypes of effective public sector CIOs. ACM Transactions on Management Information Systems (TMIS), 2(1): 5:1–5:??, March 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Du:2023:IMK

[DXC23] Kelvin Du, Frank Xing, and Erik Cambria. Incorporating multiple knowledge sources for targeted aspect-based financial sentiment analysis. ACM Transactions on Management Information Systems (TMIS), 14(3):23:1–23:??, 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3580480.

Denisenko:2024:PIM

[DZP^{+}24] Natalia Denisenko, Youzhi Zhang, Chiara Pulice, Shohini Bhattasali, Sushil Jajodia, Philip Resnik, and V. S. Subrahmanian. A psycholinguisticsinspired method to counter IP theft using fake documents. ACM Transactions on Management Information Systems (TMIS), 15(2):7:1–7:??, June 2024. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3651313.

Edla:2021:EDL

[EDBK21] Damodar Reddy Edla, Shubham Dodia, Annushree Bablani, and Venkatanareshbabu Kuppili. An efficient deep learning paradigm for deceit identification test on EEG signals. ACM Transactions on Management Information Systems (TMIS), 12(3):25:1–25:20, July 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458791.

Ermakova:2020:SPR

[EFK^{+}20] Tatiana Ermakova, Benjamin Fabian, Marta Kornacka, Scott Thiebes, and Ali Sunyaev. Security and privacy requirements for cloud computing in healthcare: Elicitation and prioritization from a patient perspective. ACM Transactions on Management Information Systems (TMIS), 11 (2):6:1–6:29, July 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386160.

Elkoumy:2022:PCP

[EFPS^{+}22] Gamal Elkoumy, Stephan A.

Fahrenkrog-Petersen, Mohammadreza Fani Sani, Agnes Koschmider, Felix Mannhardt, Saskia Nuñez Von Voigt, Majid Rafiei, and Leopold Von Waldthausen. Privacy and confidentiality in process mining: Threats and research challenges. ACM Transactions on Management Information Systems (TMIS), 13(1):11:1–11:17, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3468877.

Emami:2019:GBA

[Ema19] Hojjat Emami. A graph-based approach to person name disambiguation in Web. ACM Transactions on Management Information Systems (TMIS), 10(2):4:1–4:??, August 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3314949.

Ebrahimi:2022:QCS

[ETHJ22] Maryam Ebrahimi, Mohammad Hesam Tadayon, Mohammad Sayad Haghighi, and Alireza Jolfaei. A quantitative comparative study of data-oriented trust management schemes in Internet of Things. ACM Transactions on Management Information Systems (TMIS), 13(3):24:1–24:30, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3476248.

Edgcomb:2013:AEA

[EV13] Alex Edgcomb and Frank Vahid. Accurate and efficient algorithms that adapt to privacy-enhanced video for improved assistive monitoring.

ACM Transactions on Management Information Systems (TMIS), 4(3):14:1–14:??, October 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Eftekhari:2017:DHI

[EYS^{+}17] Saeede Eftekhari, Niam Yaraghi, Ranjit Singh, Ram D. Gopal, and R. Ramesh. Do health information exchanges deter repetition of medical services? ACM Transactions on Management Information Systems (TMIS), 8(1):2:1–2:??, May 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Fernandez:2024:ABM

[FBP24] Joaquin Delgado Fernandez, Tom Barbereau, and Orestis Papageorgiou. Agent-based model of initial token allocations: Simulating distributions post fair launch. ACM Transactions on Management Information Systems (TMIS), 15(1): 5:1–5:??, March 2024. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3649318.

Fu:2010:PPT

[FCKG10] Yu Fu, Zhiyuan Chen, Gunes Koru, and Aryya Gangopadhyay. A privacy protection technique for publishing data mining models and research data. ACM Transactions on Management Information Systems (TMIS), 1(1):7:1–7:??, December 2010. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Fridgen:2015:IBV

[FKBW15] Gilbert Fridgen, Julia Klier, Martina Beer, and Thomas Wolf. Improving business value assurance in large-scale IT projects — a quantitative method based on founded requirements assessment. ACM Transactions on Management Information Systems (TMIS), 5(3):12:1–12:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Fechner:2023:NIS

[FKK^{+}23] Pascal Fechner, Fabian König, Wolfgang Kratsch, Jannik Lockl, and Maximilian Röglinger. Near-infrared spectroscopy for bladder monitoring: a machine learning approach. ACM Transactions on Management Information Systems (TMIS), 14(2):16:1–16:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3563779.

Fan:2018:IES

[FN18] Xiangyu Fan and Xi Niu. Implementing and evaluating serendipity in delivering personalized health information. ACM Transactions on Management Information Systems (TMIS), 9(2):7:1–7:??, September 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Guo:2023:SDH

[GBK23] Tianjian Guo, Indranil Bardhan, and Anjum Khurshid. Social determinants of health and ER utilization: Role of information integration during COVID-19. ACM Transactions on Management Information Systems (TMIS), 14(4):29:1– 29:??, December 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3583077.

Gill:2013:FUM

[GH13] T. Grandon Gill and Alan R.

Hevner. A fitness-utility model for design science research. ACM Transactions on Management Information Systems (TMIS), 4(2):5:1–5:??, August 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Gopal:2020:RIU

[GHK^{+}20] Ram D. Gopal, Hooman Hidaji, Sule Nur Kutlu, Raymond A. Patterson, Erik Rolland, and Dmitry Zhdanov. Real or not?: Identifying untrustworthy news websites using third-party partnerships. ACM Transactions on Management Information Systems (TMIS), 11(3):10:1–10:20, August 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3382188.

Goes:2011:LCA

[GIYZ11] Paulo Goes, Noyan Ilk, Wei T. Yue, and J. Leon Zhao. Livechat agent assignments to heterogeneous e-customers under imperfect classification. ACM Transactions on Management Information Systems (TMIS), 2 (4):24:1–24:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Gan:2020:UDM

[GLC^{+}20] Wensheng Gan, Jerry Chun-Wei Lin, Han-Chieh Chao, Philippe Fournier-Viger, Xuan Wang, and Philip S. Yu. Utility-driven mining of trend information for intelligent system. ACM Transactions on Management Information Systems (TMIS), 11(3):14:1–14:28, August 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3391251.

Goodman:2014:BNC

[Goo14] S. E. Goodman. Building the nation’s cyber security workforce: Contributions from the CAE colleges and universities. ACM Transactions on Management Information Systems (TMIS), 5(2):6:1–6:??, July 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Gupta:2016:SCS

[GSS16] Agam Gupta, Biswatosh Saha, and Uttam K. Sarkar. Systemic concentration in sponsored search markets: The role of time window in click-through-rate computation. ACM Transactions on Management Information Systems (TMIS), 7(2): 6:1–6:??, August 2016. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Guo:2015:DPB

[GSV15] Xitong Guo, Sherry X. Sun, and Doug Vogel. A dataflow perspective for business process integration. ACM Transactions on Management Information Systems (TMIS), 5(4): 22:1–22:??, January 2015. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Gomez-Uribe:2016:NRS

[GUH16] Carlos A. Gomez-Uribe and Neil Hunt. The Netflix recommender system: Algorithms, business value, and innovation. ACM Transactions on Management Information Systems (TMIS), 6(4):13:1–13:??, January 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Gu:2021:SFR

[GVP21] Kang Gu, Soroush Vosoughi, and Temiloluwa Prioleau. SymptomID: a framework for rapid symptom identification in pandemics using news reports. ACM Transactions on Management Information Systems (TMIS), 12(4):32:1–32:17, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462441.

Gao:2022:FEH

[GYZY22] Yuan Gao, Laurence T. Yang, Yaliang Zhao, and Jing Yang. Feature extraction of highdimensional data based on J-HOSVD for cyber-physical-social systems. ACM Transactions on Management Information Systems (TMIS), 13(3): 26:1–26:21, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3483448.

Husak:2020:PCS

[HBK^{+}20] Martin Husák, Tomás Bajtos, Jaroslav Kaspar, Elias Bou-Harb, and Pavel Celeda. Predictive cyber situational awareness and personalized blacklisting: a sequential rule mining approach. ACM Transactions on Management Information Systems (TMIS), 11(4):19:1– 19:16, December 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3386250.

Huang:2014:BOT

[HBL14] Lihua Huang, Sulin Ba, and Xianghua Lu. Building online trust in a culture of Confucianism: The impact of process flexibility and perceived control. ACM Transactions on Management Information Systems (TMIS), 5(1):4:1–4:??, April 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Hu:2011:AIS

[HC11] Paul Jen-Hwa Hu and Hsinchun Chen. Analyzing information systems researchers’ productivity and impacts: a perspective on the H index. ACM Transactions on Management Information Systems (TMIS), 2(2): 7:1–7:??, June 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Hu:2012:DVP

[HCLN12] Nan Hu, Hasan Cavusoglu, Ling Liu, and Chenkai Ni. Do vendors’ pricing decisions fully reflect information in online reviews? ACM Transactions on Management Information Systems (TMIS), 3(3):16:1–16:??, October 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Hu:2023:RAH

[HDZ^{+}23] Yixiang Hu, Xiaoheng Deng, Congxu Zhu, Xuechen Chen, and Laixin Chi. Resource allocation for heterogeneous computing tasks in wirelessly powered MEC-enabled IIOT systems. ACM Transactions on Management Information Systems (TMIS), 14(1):9:1–9:??, March 2023. CODEN ????

ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571291.

Hartono:2019:WVD

[HH19] Edward Hartono and Clyde W. Holsapple. Website visual design qualities: a threefold framework. ACM Transactions on Management Information Systems (TMIS), 10 (1):1:1–1:??, May 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3309708.

Huang:2011:MTC

[HL11] Ke-Wei Huang and Zhuolun Li. A multilabel text classification algorithm for labeling risk factors in SEC form 10-K. ACM Transactions on Management Information Systems (TMIS), 2(3):18:1–18:??, October 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Huang:2021:EEM

[HL21] Peng Huang and Henry C. Lucas. Early exploration of MOOCs in the U.S. higher education: an absorptive capacity perspective. ACM Transactions on Management Information Systems (TMIS), 12(3): 22:1–22:28, July 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3456295.

Hsieh:2022:DFR

[HLCY22] Hsun-Ping Hsieh, Fandel Lin, Nai-Yu Chen, and Tzu-Hsin Yang. A decision framework to recommend cruising locations for taxi drivers under the constraint of booking information. ACM Transactions on Management Information Systems (TMIS), 13(3):27:1–27:30, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490687.

Ho:2014:SSP

[HLG14] Joyce C. Ho, Cheng H. Lee, and Joydeep Ghosh. Septic shock prediction for patients with missing data. ACM Transactions on Management Information Systems (TMIS), 5(1): 1:1–1:??, April 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

He:2022:MAC

[HLYW22] Luo He, Hongyan Liu, Yinghui Yang, and Bei Wang. A multiattention collaborative deep learning approach for blood pressure prediction. ACM Transactions on Management Information Systems (TMIS), 13(2):15:1–15:20, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3471571.

Hashmi:2016:WSN

[HMN^{+}16] Khayyam Hashmi, Zaki Malik, Erfan Najmi, Amal Alhosban, and Brahim Medjahed. A Web service negotiation management and QoS dependency modeling framework. ACM Transactions on Management Information Systems (TMIS), 7 (2):5:1–5:??, August 2016. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Hevner:2023:RCD

[HS23] Alan Hevner and Veda Storey. Research challenges for the design of human-artificial intelligence systems (HAIS). ACM Transactions on Management Information Systems (TMIS), 14(1):10:1–10:??, March 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3549547.

Ha:2017:PPE

[HSK17] Tuan Minh Ha, Masaki Samejima, and Norihisa Komoda. Power and performance estimation for fine-grained server power capping via controlling heterogeneous applications. ACM Transactions on Management Information Systems (TMIS), 8(4):11:1–11:??, September 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Han:2019:CTR

[HYG19] Xu Han, Niam Yaraghi, and Ram Gopal. Catching them red-handed: Optimizing the nursing homes’ rating system.

ACM Transactions on Management Information Systems (TMIS), 10(2):7:1–7:??, August 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3325522.

Huang:2022:IMA

[HYYY22] Shi Ming Huang, David C. Yen, Ting Jyun Yan, and Yi Ting Yang. An intelligent mechanism to automatically discover emerging technology trends: Exploring regulatory technology. ACM Transactions on Management Information Systems (TMIS), 13(2):17:1–17:29, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485187.

Huang:2012:TNP

[HZZ12] Zan Huang, Huimin Zhao, and Dan Zhu. Two new predictiondriven approaches to discrete choice prediction. ACM Transactions on Management Information Systems (TMIS), 3(2): 9:1–9:??, July 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Jiang:2015:CCO

[JADT15] Jie Jiang, Huib Aldewereld, Virginia Dignum, and Yao-Hua Tan. Compliance checking of organizational interactions. ACM Transactions on Management Information Systems (TMIS), 5(4):23:1–23:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Jureta:2015:RPA

[JBEM15] Ivan J. Jureta, Alexander Borgida, Neil A. Ernst, and John Mylopoulos. The requirements problem for adaptive systems. ACM Transactions on Management Information Systems (TMIS), 5(3):17:1–17:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Jain:2015:SBS

[JCMR15] Radhika Jain, Lan Cao, Kannan Mohan, and Balasubramaniam Ramesh. Situated boundary spanning: an empirical investigation of requirements engineering practices in product family development. ACM Transactions on Management Information Systems (TMIS), 5(3):16:1–16:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Janiesch:2023:AUP

[JFI^{+}23] Christian Janiesch, Marcus Fischer, Florian Imgrund, Adrian Hofmann, and Axel Winkelmann. An architecture using payment channel networks for blockchain-based Wi-Fi sharing. ACM Transactions on Management Information Systems (TMIS), 14(1):1:1–1:??, March 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529097.

Jiang:2019:MAI

[JHC19] Jian-Min Jiang, Zhong Hong, and Yangyang Chen. Modeling and analyzing incremental natures of developing software. ACM Transactions on Management Information Systems (TMIS), 10(2): 5:1–5:??, August 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3333535.

Jannach:2019:MBV

[JJ19] Dietmar Jannach and Michael Jugovac. Measuring the business value of recommender systems. ACM Transactions on Management Information Systems (TMIS), 10(4):16:1– 16:??, December 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3370082.

Jarke:2015:ECS

[JL15] Matthias Jarke and Kalle Lyytinen. Editorial: “Complexity of Systems Evolution: Requirements Engineering Perspective”. ACM Transactions on Management Information Systems (TMIS), 5(3):11:1– 11:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Joshi:2023:LNL

[JNDZ23] Amogh Manoj Joshi, Deepak Ranjan Nayak, Dibyasundar Das, and Yudong Zhang. LiMS-Net: a lightweight multi-scale CNN for COVID-19 detection from chest CT scans. ACM Transactions on Management Information Systems (TMIS), 14(1): 5:1–5:??, March 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3551647.

Jain:2019:ISS

[JRYY19] Hemant Jain, T. S. Raghu, Victoria Yoon, and Wei Thoo Yue. Introduction to special section based on papers presented at the Workshop on Information Technology and Systems, 2017. ACM Transactions on Management Information Systems (TMIS), 10(2): 6:1–6:??, August 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3342557.

Kakar:2017:IRB

[Kak17] Adarsh Kumar Kakar. Investigating the relationships between the use contexts, user perceived values, and loyalty to a software product. ACM Transactions on Management Information Systems (TMIS), 8 (1):3:1–3:??, May 2017. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Kane:2011:MSI

[Kan11] Gerald C. Kane. A multimethod study of information quality in wiki collaboration. ACM Transactions on Management Information Systems (TMIS), 2(1):4:1–4:??, March 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Khosla:2013:ECM

[KC13] Rajiv Khosla and Mei-Tai Chu.

Embodying care in Matilda: an affective communication robot for emotional wellbeing of older people in Australian residential care facilities. ACM Transactions on Management Information Systems (TMIS), 4 (4):18:1–18:??, December 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Ketter:2018:ISS

[KCSTM18] Wolfgang Ketter, John Collins, Maytal Saar-Tsechansky, and Ori Marom. Information systems for a smart electricity grid: Emerging challenges and opportunities. ACM Transactions on Management Information Systems (TMIS), 9 (3):10:1–10:??, November 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3230712.

Kratzwald:2019:PQA

[KF19] Bernhard Kratzwald and Stefan Feuerriegel. Putting question-answering systems into practice: Transfer learning for efficient domain customization. ACM Transactions on Management Information Systems (TMIS), 9(4): 15:1–15:??, March 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3309706.

Killoran:2023:IIC

[KJM23] Jayson Andrew Killoran, Tracy A. Jenkin, and Jasmin Manseau. ICT interactions and COVID-19 — a theorization across two pandemic waves. ACM Transactions on Management Information Systems (TMIS), 14 (4):30:1–30:??, December 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3597938.

Krishnamurthy:2016:PDP

[KJRD16] Rajiv Krishnamurthy, Varghese Jacob, Suresh Radhakrishnan, and Kutsal Dogan. Peripheral developer participation in open source projects: an empirical analysis. ACM Transactions on Management Information Systems (TMIS), 6(4): 14:1–14:??, January 2016. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Kumar:2022:DBP

[KKB^{+}22] Ankit Kumar, Abhishek Kumar, Ali Kashif Bashir, Mamoon Rashid, V. D. Ambeth Kumar, and Rupak Kharel. Distance based pattern driven mining for outlier detection in high dimensional big dataset. ACM Transactions on Management Information Systems (TMIS), 13(1): 8:1–8:17, March 2022. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3469891.

Kartal:2019:DPV

[KLL19] Hasan B. Kartal, Xiaoping Liu, and Xiao-Bai Li. Differential privacy for the vast majority. ACM Transactions on Management Information Systems (TMIS), 10(2): 8:1–8:??, August 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3329717.

Kucklick:2023:TAI

[KM23] Jan-Peter Kucklick and Oliver Müller. Tackling the accuracyinterpretability trade-off: Interpretable deep learning models for satellite image-based real estate appraisal. ACM Transactions on Management Information Systems (TMIS), 14(1): 6:1–6:??, March 2023. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3567430.

Kharkwal:2021:UOD

[KOH^{+}21] Himanshu Kharkwal, Dakota Olson, Jiali Huang, Abhiraj Mohan, Ankur Mani, and Jaideep Srivastava. University operations during a pandemic: a flexible decision analysis toolkit. ACM Transactions on Management Information Systems (TMIS), 12(4):35:1– 35:24, December 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3460125.

Kasiri:2013:ROS

[KS13] Narges Kasiri and Ramesh Sharda. Real options and system dynamics for information technology investment decisions: Application to RFID adoption in retail. ACM Transactions on Management Information Systems (TMIS), 4(3): 11:1–11:??, October 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

King:2015:CCR

[KS15] John Leslie King and Carl P. Simon. Complications with complexity in requirements. ACM Transactions on Management Information Systems (TMIS), 5(3):13:1–13:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Kostkova:2014:SUT

[KSS14] Patty Kostkova, Martin Szomszor, and Connie St. Luis. #swineflu: The use of Twitter as an early warning and risk communication tool in the 2009 swine flu pandemic. ACM Transactions on Management Information Systems (TMIS), 5 (2):8:1–8:??, July 2014. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Keller:2015:CMR

[KTF15] Thorben Keller, Frédéric Thiesse, and Elgar Fleisch. Classification models for RFID-based real-time detection of process events in the supply chain: an empirical study. ACM Transactions on Management Information Systems (TMIS), 5(4): 25:1–25:??, January 2015. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Kul:2021:ACI

[KUH21] Gökhan Kul, Shambhu Upadhyaya, and Andrew Hughes. An analysis of complexity of insider attacks to databases. ACM Transactions on Management Information Systems (TMIS), 12(1):4:1–4:18, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3391231.

Kuo:2011:LAG

[KY11] Feng-Yang Kuo and Chun-Po Yin. A linguistic analysis of group support systems interactions for uncovering social realities of organizations. ACM Transactions on Management Information Systems (TMIS), 2 (1):3:1–3:??, March 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Kang:2019:HAO

[KZ19] Yin Kang and Lina Zhou. Helpfulness assessment of online reviews: The role of semantic hierarchy of product features. ACM Transactions on Management Information Systems (TMIS), 10(3):12:1– 12:??, November 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3365538.

Kesan:2020:ACI

[KZ20] Jay P. Kesan and Linfeng Zhang. Analysis of cyber incident categories based on losses. ACM Transactions on Management Information Systems (TMIS), 11(4):25:1–25:28, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3418288.

Lisetti:2013:CHY

[LAYR13] Christine Lisetti, Reza Amini, Ugan Yasavur, and Naphtali Rishe. I can help you change! An empathic virtual agent delivers behavior change health interventions. ACM Transactions on Management Information Systems (TMIS), 4 (4):19:1–19:??, December 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lo:2019:PWT

[LC19] Kar Kei Lo and Michael Chau. A penny is worth a thousand? Investigating the relationship between social media and penny stocks. ACM Transactions on Management Information Systems (TMIS), 9(4): 14:1–14:??, March 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3309704.

Lim:2013:BIA

[LCC13] Ee-Peng Lim, Hsinchun Chen, and Guoqing Chen. Business intelligence and analytics: Research directions. ACM Transactions on Management Information Systems (TMIS), 3(4): 17:1–17:??, January 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lu:2017:SLE

[LCC17] Yan Lu, Michael Chau, and Patrick Y. K. Chau. Are sponsored links effective? Investigating the impact of trust in search engine advertising. ACM Transactions on Management Information Systems (TMIS), 7(4):12:1–12:??, January 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lv:2022:CIS

[LCF^{+}22] Zhihan Lv, Dongliang Chen, Hailin Feng, Amit Kumar Singh, Wei Wei, and Haibin Lv. Computational intelligence in security of digital twins big graphic data in cyberphysical systems of smart cities. ACM Transactions on Management Information Systems (TMIS), 13(4):39:1–39:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3522760.

Lu:2020:AAW

[LCS^{+}20] Haibing Lu, Xi Chen, Junmin Shi, Jaideep Vaidya, Vijayalakshmi Atluri, Yuan Hong, and Wei Huang. Algorithms and applications to weighted rank-one binary matrix factorization. ACM Transactions on Management Information Systems (TMIS), 11 (2):7:1–7:33, July 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386599.

Li:2024:EFF

[LCWH24] Pei-Xuan Li, Yu-En Chang, Ming-Chun Wei, and Hsun-Ping Hsieh. Estimating future financial development of urban areas for deploying bank branches: a local-regional interpretable model. ACM Transactions on Management Information Systems (TMIS), 15(2):8:1–8:??, June 2024. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3656479.

Li:2018:SLR

[LFS18] Zhepeng (Lionel) Li, Xiao Fang, and Olivia R. Liu Sheng.

A survey of link recommendation for social networks: Methods, theoretical foundations, and future research directions. ACM Transactions on Management Information Systems (TMIS), 9(1):1:1–1:??, February 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lin:2022:SSJ

[LGW^{+}22] Qi Lin, Wensheng Gan, Yongdong Wu, Jiahui Chen, and Chien-Ming Chen. Smart system: Joint utility and frequency for pattern classification. ACM Transactions on Management Information Systems (TMIS), 13(4):43:1–43:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531480.

Lee:2015:OBM

[LHT15] Yen-Hsien Lee, Paul Jen-Hwa Hu, and Ching-Yi Tu. Ontology-based mapping for automated document management: a concept-based technique for word mismatch and ambiguity problems in document clustering. ACM Transactions on Management Information Systems (TMIS), 6(1): 4:1–4:??, March 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Li:2023:ISI

[LKPZ23] Xin Li, Juhee Kwon, Balaji Padmanabhan, and Pengzhu Zhang. Introduction to the special issue on IT-enabled business management and decision making in the (post) Covid-19 era. ACM Transactions on Management Information Systems (TMIS), 14(4):28:1–28:??, December 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3627995.

Li:2015:NBB

[LKZ^{+}15] Shing-Han Li, Yu-Cheng Kao, Zong-Cyuan Zhang, Ying-Ping Chuang, and David C. Yen. A network behavior-based Botnet detection mechanism using PSO and K-means. ACM Transactions on Management Information Systems (TMIS), 6(1):3:1–3:??, February 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lo:2024:NMG

[LL24] Pei-Chi Lo and Ee-Peng Lim. Non-monotonic generation of knowledge paths for context understanding. ACM Transactions on Management Information Systems (TMIS), 15(1): 1:1–1:??, March 2024. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3627994.

Lv:2022:NML

[LLF^{+}22] Zhihan Lv, Ranran Lou, Hailin Feng, Dongliang Chen, and Haibin Lv. Novel machine learning for big data analytics in intelligent support information management systems. ACM Transactions on Management Information Systems (TMIS), 13(1):7:1–7:21, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469890.

Lau:2011:TMP

[LLK^{+}11] Raymond Y. K. Lau, S. Y.

Liao, Ron Chi-Wai Kwok, Kaiquan Xu, Yunqing Xia, and Yuefeng Li. Text mining and probabilistic language modeling for online review spam detection. ACM Transactions on Management Information Systems (TMIS), 2(4):25:1–25:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lin:2011:SPM

[LLKC11] Ming-Chih Lin, Anthony J. T.

Lee, Rung-Tai Kao, and Kuo-Tay Chen. Stock price movement prediction using representative prototypes of financial reports. ACM Transactions on Management Information Systems (TMIS), 2(3):19:1–19:??, October 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lui:2016:EMC

[LP16] Tsz-Wai Lui and Gabriele Piccoli. The effect of a multichannel customer service system on customer service and financial performance. ACM Transactions on Management Information Systems (TMIS), 7(1):2:1– 2:??, March 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Liu:2011:WDW

[LR11] Jun Liu and Sudha Ram. Who does what: Collaboration patterns in the Wikipedia and their impact on article quality. ACM Transactions on Management Information Systems (TMIS), 2(2):11:1–11:??, June 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lukyanenko:2017:ACC

[LS17] Roman Lukyanenko and Binny M. Samuel. Are all classes created equal? Increasing precision of conceptual modeling grammars. ACM Transactions on Management Information Systems (TMIS), 8(4):14:1–14:??, September 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Li:2021:CSS

[LSM^{+}21] Junye Li, Aryan Sharma, Deepak Mishra, Gustavo Batista, and Aruna Seneviratne. COVIDsafe spatial occupancy monitoring using OFDM-based features and passive WiFi samples. ACM Transactions on Management Information Systems (TMIS), 12(4):34:1–34:24, December 2021. CODEN ????

ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472668.

Liu:2015:WNS

[LSS15] Dengpan Liu, Sumit Sarkar, and Chelliah Sriskandarajah. Who’s next? Scheduling personalization services with variable service times. ACM Transactions on Management Information Systems (TMIS), 6(2): 8:1–8:??, July 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Li:2020:IIE

[LSS20] Hongfei Li, Ramesh Shankar, and Jan Stallaert. Invested or indebted: Ex-ante and ex-post reciprocity in online knowledge sharing communities. ACM Transactions on Management Information Systems (TMIS), 11(1):1:1–1:26, April 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3371388.

Lin:2022:ISI

[LSSD22] Jerry Chun-Wei Lin, Nachiketa Sahoo, Gautam Srivastava, and Weiping Ding. Introduction to the special issue on patterndriven mining, analytics, and prediction for decision making, Part II. ACM Transactions on Management Information Systems (TMIS), 13(3):23:1–23:3, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3512468.

Li:2022:IBS

[LST22] Guangrui (Kayla) Li, Mike K. P. So, and Kar Yan Tam. Identifying the big shots — a quantile-matching way in the big data context. ACM Transactions on Management Information Systems (TMIS), 13(2): 19:1–19:30, June 2022. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3490395.

Liaskos:2016:SRO

[LT16] Christos Liaskos and Ageliki Tsioliaridou. Service ratiooptimal, content coherenceaware data push systems. ACM Transactions on Management Information Systems (TMIS), 6(4):15:1–15:??, January 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lu:2012:CRC

[LTC^{+}12] Hsin-Min Lu, Feng-Tse Tsai, Hsinchun Chen, Mao-Wei Hung, and Shu-Hsing Li. Credit rating change modeling using news and financial ratios. ACM Transactions on Management Information Systems (TMIS), 3(3):14:1–14:??, October 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Leroy:2023:ISI

[LTL23] Gondy Leroy, Bengisu Tulu, and Xiao Liu. Introduction to the special issue on design and data science research in healthcare. ACM Transactions on Management Information Systems (TMIS), 14(2):12:1– 12:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579646.

Li:2016:RTA

[LYC16] Shing-Han Li, David C. Yen, and Ying-Ping Chuang. A realtime audit mechanism based on the compression technique. ACM Transactions on Management Information Systems (TMIS), 7(2):4:1–4:??, August 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Leng:2021:LIC

[LZK21] Yan Leng, Jinhua Zhao, and Haris Koutsopoulos. Leveraging individual and collective regularity to profile and segment user locations from mobile phone data. ACM Transactions on Management Information Systems (TMIS), 12(3): 20:1–20:22, July 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3449042.

Liang:2021:USM

[LZLL21] Guanqing Liang, Jingxin Zhao, Helena Yan Ping Lau, and Cane Wing-Ki Leung. Using social media to analyze public concerns and policy responses to COVID-19 in Hong Kong. ACM Transactions on Management Information Systems (TMIS), 12(4):30:1–30:20, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460124.

Lai:2019:TVK

[LZW^{+}19] Jianwei Lai, Dongsong Zhang, Sen Wang, Isil Doga Yakut Kilic, and Lina Zhou. Thumb-Stroke: a virtual keyboard in support of sight-free and onehanded text entry on touchscreen mobile devices. ACM Transactions on Management Information Systems (TMIS), 10(3):11:1–11:??, November 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3343858.

Lu:2016:CCB

[LZX16] Xianghua Lu, Xia Zhao, and Ling Xue. Is combining contextual and behavioral targeting strategies effective in online advertising? ACM Transactions on Management Information Systems (TMIS), 7(1):1:1– 1:??, March 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Liu:2022:TSC

[LZX^{+}22a] Genggeng Liu, Ruping Zhou, Saijuan Xu, Yuhan Zhu, Wenzhong Guo, Yeh-Cheng Chen, and Guolong Chen. Two-stage competitive particle swarm optimization based timingdriven X-routing for IC design under smart manufacturing. ACM Transactions on Management Information Systems (TMIS), 13(4):41:1–41:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531328.

Liu:2022:PDX

[LZX^{+}22b] Genggeng Liu, Yuhan Zhu, Saijuan Xu, Hao Tang, and Yeh-Cheng Chen. Performancedriven X-architecture routing algorithm for artificial intelligence chip design in smart manufacturing. ACM Transactions on Management Information Systems (TMIS), 13 (4):38:1–38:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3519422.

Liu:2022:EFP

[LZZ^{+}22] Xin Liu, Liang Zheng, Weishan Zhang, Jiehan Zhou, Shuai Cao, and Shaowen Yu. An evolutive frequent pattern treebased incremental knowledge discovery algorithm. ACM Transactions on Management Information Systems (TMIS), 13(3):30:1–30:20, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495213.

Masud:2011:CBM

[MAKH^{+}11] Mohammad M. Masud, Tahseen M. Al-Khateeb, Kevin W. Hamlen, Jing Gao, Latifur Khan, Jiawei Han, and Bhavani Thuraisingham. Cloud-based malware detection for evolving data streams. ACM Transactions on Management Information Systems (TMIS), 2(3): 16:1–16:??, October 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Marshall:2022:MLS

[MCCC22] Byron Marshall, Michael Curry, Robert E. Crossler, and John Correia. Machine learning and survey-based predictors of InfoSec non-compliance. ACM Transactions on Management Information Systems (TMIS), 13(2):13:1–13:20, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3466689.

Mcknight:2011:TST

[MCTC11] D. Harrison Mcknight, Michelle Carter, Jason Bennett Thatcher, and Paul F. Clay. Trust in a specific technology: an investigation of its components and measures. ACM Transactions on Management Information Systems (TMIS), 2(2): 12:1–12:??, June 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Mannino:2017:DES

[MFBK^{+}17] Michael Mannino, Joel Fredrickson, Farnoush Banaei-Kashani, Iris Linck, and Raghda Alqurashi Raghda. Development and evaluation of a similarity measure for medical event sequences. ACM Transactions on Management Information Systems (TMIS), 8(2–3):8:1–8:??, August 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Mirani:2013:BBI

[MH13] Rajesh Mirani and Anju Harpalani. Business benefits or incentive maximization? impacts of the Medicare EHR incentive program at acute care hospitals. ACM Transactions on Management Information Systems (TMIS), 4(4):20:1– 20:??, December 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Ma:2022:SME

[MHC^{+}22] Wanlun Ma, Xiangyu Hu, Chao Chen, Sheng Wen, Kkwang Raymond Choo, and Yang Xiang. Social media event prediction using DNN with feedback mechanism. ACM Transactions on Management Information Systems (TMIS), 13(3):33:1–33:24, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522759.

Mohammadi:2019:SAB

[MHT19] Majid Mohammadi, Wout Hofman, and Yao-Hua Tan. Simulated annealing-based ontology matching. ACM Transactions on Management Information Systems (TMIS), 10 (1):3:1–3:??, May 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3314948.

Malhotra:2012:HVT

[MM12] Arvind Malhotra and Ann Majchrzak. How virtual teams use their virtual workspace to coordinate knowledge. ACM Transactions on Management Information Systems (TMIS), 3(1): 6:1–6:??, April 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Marx:2011:SPR

[MMW11] Frederik Marx, Jörg H. Mayer, and Robert Winter. Six principles for redesigning executive information systems-findings of a survey and evaluation of a prototype. ACM Transactions on Management Information Systems (TMIS), 2 (4):26:1–26:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Mathew:2013:DPP

[MO13] George Mathew and Zoran Obradovic. Distributed privacy-preserving decision support system for highly imbalanced clinical data. ACM Transactions on Management Information Systems (TMIS), 4(3):12:1–12:??, October 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Miksa:2022:ARD

[MOR22] Tomasz Miksa, Simon Oblasser, and Andreas Rauber. Automating research data management using machine-actionable data management plans. ACM Transactions on Management Information Systems (TMIS), 13(2):18:1–18:22, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3490396.

Mangino:2020:ISI

[MPBH20] Antonio Mangino, Morteza Safaei Pour, and Elias Bou-Harb. Internet-scale insecurity of consumer Internet of Things: an empirical measurements perspective. ACM Transactions on Management Information Systems (TMIS), 11(4):21:1–21:24, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3394504.

Mishra:2023:EED

[MRKE23] Rahul Mishra, Dharavath Ramesh, Salil S. Kanhere, and Damodar Reddy Edla. Enabling efficient deduplication and secure decentralized public auditing for cloud storage: a redactable blockchain approach. ACM Transactions on Management Information Systems (TMIS), 14(3):21:1– 21:??, 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3578555.

Mudgerikar:2020:EBI

[MSB20] Anand Mudgerikar, Puneet Sharma, and Elisa Bertino. Edge-based intrusion detection for IoT devices. ACM Transactions on Management Information Systems (TMIS), 11 (4):18:1–18:21, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3382159.

Mukherjee:2017:ARB

[MSD17] Anik Mukherjee, R. P. Sundarraj, and Kaushik Dutta. Apriori rule-based in-app ad selection online algorithm for improving supply-side platform revenues. ACM Transactions on Management Information Systems (TMIS), 8(2–3):10:1– 10:??, August 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Morid:2023:TSP

[MSD23] Mohammad Amin Morid, Olivia R. Liu Sheng, and Joseph Dunbar. Time series prediction using deep learning methods in healthcare. ACM Transactions on Management Information Systems (TMIS), 14(1):2:1–2:??, March 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3531326.

Mehrotra:2020:PPD

[MSU^{+}20] Sharad Mehrotra, Shantanu Sharma, Jeffrey D. Ullman, Dhrubajyoti Ghosh, Peeyush Gupta, and Anurag Mishra. PANDA: Partitioned data security on outsourced sensitive and non-sensitive data. ACM Transactions on Management Information Systems (TMIS), 11(4):23:1–23:41, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3397521.

Mendling:2018:BBP

[MWV^{+}18] Jan Mendling, Ingo Weber, Wil Van Der Aalst, Jan Vom Brocke and Cristina Cabanillas, Florian Daniel, Søren Debois, Claudio Di Ciccio, Marlon Dumas, Schahram Dustdar, Avigdor Gal, Luciano García-Bañuelos, Guido Governatori, Richard Hull, Marcello La Rosa, Henrik Leopold, Frank Leymann, Jan Recker, Manfred Reichert, Hajo A. Reijers, Stefanie Rinderle-Ma, Andreas Solti, Michael Rosemann, Stefan Schulte, Munindar P. Singh, Tijs Slaats, Mark Staples, Barbara Weber, Matthias Weidlich, Mathias Weske, Xiwei Xu, and Liming Zhu. Blockchains for business process management — challenges and opportunities.

ACM Transactions on Management Information Systems (TMIS), 9(1):4:1–4:??, February 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Nunamaker:2011:TBV

[NB11] Jay F. Nunamaker, Jr. and Robert O. Briggs. Toward a broader vision for Information Systems. ACM Transactions on Management Information Systems (TMIS), 2(4):20:1–20:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Nawaz:2022:MHU

[NFVY^{+}22] M. Saqib Nawaz, Philippe Fournier-Viger, Unil Yun, Youxi Wu, and Wei Song. Mining high utility itemsets with Hill climbing and simulated annealing. ACM Transactions on Management Information Systems (TMIS), 13(1):4:1–4:22, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3462636.

Nong:2022:ARC

[NHL22] Mengxin Nong, Lingfeng Huang, and Mingtao Liu. Allocation of resources for cloud survivability in smart manufacturing. ACM Transactions on Management Information Systems (TMIS), 13(4):45:1–45:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3533701.

Niu:2022:CMT

[NJC^{+}22] Shuteng Niu, Yushan Jiang, Bowen Chen, Jian Wang, Yongxin Liu, and Houbing Song. Cross-modality transfer learning for image-text information management. ACM Transactions on Management Information Systems (TMIS), 13(1):5:1–5:14, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3464324.

Ni:2020:MLD

[NLLZ20] Li Ni, Wenjian Luo, Nannan Lu, and Wenjie Zhu. Mining the local dependency itemset in a products network. ACM Transactions on Management Information Systems (TMIS), 11(1):3:1–3:31, April 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3384473.

Niederman:2012:DSA

[NM12] Fred Niederman and Salvatore T. March. Design science and the accumulation of knowledge in the information systems discipline. ACM Transactions on Management Information Systems (TMIS), 3(1): 1:1–1:??, April 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Nekvi:2015:IRC

[NM15] Md Rashed I. Nekvi and Nazim H. Madhavji. Impediments to regulatory compliance of requirements in contractual systems engineering projects: a case study. ACM Transactions on Management Information Systems (TMIS), 5(3): 15:1–15:??, January 2015. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Nussbaumer:2012:EVC

[NMS12] Philipp Nussbaumer, Inu Matter, and Gerhard Schwabe. “enforced” vs. “casual” transparency — findings from ITsupported financial advisory encounters. ACM Transactions on Management Information Systems (TMIS), 3(2): 11:1–11:??, July 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Nakashima:2021:AFS

[NSK^{+}21] Makiya Nakashima, Alex Sim, Youngsoo Kim, Jonghyun Kim, and Jinoh Kim. Automated feature selection for anomaly detection in network traffic data. ACM Transactions on Management Information Systems (TMIS), 12(3):18:1–18:28, July 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3446636.

Ng:2021:LSM

[NST21] Ka Chung Ng, Mike K. P. So, and Kar Yan Tam. A latent space modeling approach to interfirm relationship analysis. ACM Transactions on Management Information Systems (TMIS), 12(2):10:1–10:44, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424240.

Ngo-Ye:2012:AOR

[NYS12] Thomas L. Ngo-Ye and Atish P.

Sinha. Analyzing online review helpfulness using a regressional ReliefF-enhanced text mining method. ACM Transactions on Management Information Systems (TMIS), 3(2):10:1–10:??, July 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Ouyang:2021:DRS

[OAHY21] Chun Ouyang, Michael Adams, Arthur H. M. Ter Hofstede, and Yang Yu. Design and realisation of scalable business process management systems for deployment in the cloud. ACM Transactions on Management Information Systems (TMIS), 12(4):36:1–36:26, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460123.

Orman:2013:BIT

[Orm13] Levent V. Orman. Bayesian inference in trust networks. ACM Transactions on Management Information Systems (TMIS), 4 (2):7:1–7:??, August 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Onumo:2021:AME

[OUAC21] Aristotle Onumo, Irfan Ullah-Awan, and Andrea Cullen. Assessing the moderating effect of security technologies on employees compliance with cybersecurity control procedures. ACM Transactions on Management Information Systems (TMIS), 12(2):11:1–11:29, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3424282.

Purao:2018:MLC

[PBT18] Sandeep Purao, Narasimha Bolloju, and Chuan-Hoo Tan. A modeling language for conceptual design of systems integration solutions. ACM Transactions on Management Information Systems (TMIS), 9(2): 8:1–8:??, September 2018. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Parameshwarappa:2021:ADA

[PCK21] Pooja Parameshwarappa, Zhiyuan Chen, and Günes Koru. Anonymization of daily activity data by using l-diversity privacy model. ACM Transactions on Management Information Systems (TMIS), 12(3):23:1– 23:21, July 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3456876.

Park:2019:ECR

[PCLL19] Jiyong Park, Daegon Cho, Jae Kyu Lee, and Byungtae Lee. The economics of cybercrime: The role of broadband and socioeconomic status. ACM Transactions on Management Information Systems (TMIS), 10(4):13:1– 13:??, December 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3351159.

Pervin:2013:FSC

[PFD^{+}13] Nargis Pervin, Fang Fang, Anindya Datta, Kaushik Dutta, and Debra Vandermeer. Fast, scalable, and context-sensitive detection of trending topics in microblog post streams. ACM Transactions on Management Information Systems (TMIS), 3(4):19:1–19:??, January 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Pal:2019:REC

[PGPB19] Ranjan Pal, Leana Golubchik, Konstantions Psounis, and Tathagata Bandyopadhyay. On robust estimates of correlated risk in cyber-insured IT firms: a first look at optimal AIbased estimates under “small” data. ACM Transactions on Management Information Systems (TMIS), 10(3):9:1– 9:??, November 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3351158.

Parameswarath:2022:UEP

[PGS22] Rohini Poolat Parameswarath, Prosanta Gope, and Biplab Sikdar. User-empowered privacypreserving authentication protocol for electric vehicle charging based on decentralized identity and verifiable credential. ACM Transactions on Management Information Systems (TMIS), 13(4):44:1–44:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532869.

Padmanabhan:2011:IOS

[PHCS11] Balaji Padmanabhan, Alan Hevner, Michael Cuenco, and Crystal Shi. From information to operations: Service quality and customer retention. ACM Transactions on Management Information Systems (TMIS), 2 (4):21:1–21:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Pal:2021:WCC

[PHL^{+}21] Ranjan Pal, Ziyuan Huang, Sergey Lototsky, Xinlong Yin, Mingyan Liu, Jon Crowcroft, Nishanth Sastry, Swades De, and Bodhibrata Nag. Will catastrophic cyber-risk aggregation thrive in the IoT age? A cautionary economics tale for (re-)insurers and likes. ACM Transactions on Management Information Systems (TMIS), 12(2):17:1–17:36, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3446635.

Pika:2017:MRP

[PLW^{+}17] Anastasiia Pika, Michael Leyer, Moe T. Wynn, Colin J. Fidge, Arthur H. M. Ter Hofstede, and Wil M. P. Van Der Aalst. Mining resource profiles from event logs. ACM Transactions on Management Information Systems (TMIS), 8(1):1:1– 1:??, May 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Pierazzi:2020:DDC

[PMH^{+}20] Fabio Pierazzi, Ghita Mezzour, Qian Han, Michele Colajanni, and V. S. Subrahmanian. A data-driven characterization of modern Android spyware. ACM Transactions on Management Information Systems (TMIS), 11 (1):4:1–4:38, April 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3382158.

Pika:2022:CBP

[POtH22] Anastasiia Pika, Chun Ouyang, and Arthur H. M. ter Hofstede. Configurable batchprocessing discovery from event logs. ACM Transactions on Management Information Systems (TMIS), 13(3):28:1–28:25, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490394.

Pal:2020:WCB

[PPC^{+}20] Ranjan Pal, Konstantinos Psounis, Jon Crowcroft, Frank Kelly, Pan Hui, Sasu Tarkoma, Abhishek Kumar, John Kelly, Aritra Chatterjee, Leana Golubchik, Nishanth Sastry, and Bodhibrata Nag. When are cyber blackouts in modern service networks likely?: a network oblivious theory on cyber (re)insurance feasibility. ACM Transactions on Management Information Systems (TMIS), 11(2):5:1–5:38, July 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386159.

Pilato:2023:MSS

[PPG^{+}23] Giovanni Pilato, Fabio Persia, Mouzhi Ge, Theodoros Chondrogiannis, and Daniela D’Auria. A modular social sensing system for personalized orienteering in the COVID-19 era. ACM Transactions on Management Information Systems (TMIS), 14(4):31:1–31:??, December 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3615359.

Partington:2015:PMC

[PWS^{+}15] Andrew Partington, Moe Wynn, Suriadi Suriadi, Chun Ouyang, and Jonathan Karnon. Process mining for clinical processes: a comparative analysis of four Australian hospitals. ACM Transactions on Management Information Systems (TMIS), 5(4):19:1–19:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Peng:2011:LSC

[PZH11] Jing Peng, Daniel D. Zeng, and Zan Huang. Latent subjectcentered modeling of collaborative tagging: an application in social search. ACM Transactions on Management Information Systems (TMIS), 2(3): 15:1–15:??, October 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Pytel:2024:DDT

[PZW24] Norman Pytel, Christian Ziegler, and Axel Winkelmann. From dissonance to dialogue: a token-based approach to bridge the gap between manufacturers and customers. ACM Transactions on Management Information Systems (TMIS), 15(1): 3:1–3:??, March 2024. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3639058.

Qiu:2021:MDP

[QGRT21] Lin Qiu, Sruthi Gorantla, Vaibhav Rajan, and Bernard C. Y. Tan. Multi-disease predictive analytics: a clinical knowledgeaware approach. ACM Transactions on Management Information Systems (TMIS), 12(3): 19:1–19:34, July 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3447942.

Quattrocchi:2024:DPS

[QvdHT24] Giovanni Quattrocchi, Willem-Jan van den Heuvel, and Damian Andrew Tamburri. The data product-service composition frontier: a hybrid learning approach. ACM Transactions on Management Information Systems (TMIS), 15(1):6:1–6:??, March 2024. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3649319.

Robinson:2012:DDB

[RADS12] William N. Robinson, Arash Akhlaghi, Tianjie Deng, and Ali Raza Syed. Discovery and diagnosis of behavioral transitions in patient event streams. ACM Transactions on Management Information Systems (TMIS), 3(1):4:1–4:??, April 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Rashid:2022:ADC

[RASHD22] A. N. M. Bazlur Rashid, Mohiuddin Ahmed, Leslie F. Sikos, and Paul Haskell-Dowland. Anomaly detection in cybersecurity datasets via cooperative co-evolution-based feature selection. ACM Transactions on Management Information Systems (TMIS), 13(3):29:1–29:39, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495165.

Rubin:2023:UTA

[RB23] Eran Rubin and Izak Benbasat.

Using Toulmin’s argumentation model to enhance trust in analytics-based advice giving systems. ACM Transactions on Management Information Systems (TMIS), 14(3): 22:1–22:??, 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580479.

Russo:2018:MMI

[RCFT18] Daniel Russo, Paolo Ciancarini, Tommaso Falasconi, and Massimo Tomasi. A meta-model for information systems quality: a mixed study of the financial sector. ACM Transactions on Management Information Systems (TMIS), 9 (3):11:1–11:??, November 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3230713.

Romero:2022:IDP

[RCM^{+}22] Esteban Elias Romero, Carlos David Camacho, Carlos Enrique Montenegro, Óscar Esneider Acosta, Rubén González Crespo, Elvis Eduardo Gaona, and Marcelo Herrera Martínez. Integration of DevOps practices on a noise monitor system with CircleCI and Terraform. ACM Transactions on Management Information Systems (TMIS), 13(4):36:1–36:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505228.

Ren:2022:AMD

[RCW22] Bin Ren, Yuquiang Chen, and Fujie Wang. Application massive data processing platform for smart manufacturing based on optimization of data storage. ACM Transactions on Management Information Systems (TMIS), 13(4):37:1–37:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3508395.

Rezvani:2020:RRS

[RR20] Mohsen Rezvani and Mojtaba Rezvani. A randomized reputation system in the presence of unfair ratings. ACM Transactions on Management Information Systems (TMIS), 11 (1):2:1–2:16, April 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3384472.

Ranathunga:2021:MRM

[RRN21] Dinesha Ranathunga, Matthew Roughan, and Hung Nguyen. Mathematical reconciliation of medical privacy policies. ACM Transactions on Management Information Systems (TMIS), 12(1):5:1–5:18, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3397520.

Roy:2015:MOU

[RSM^{+}15] Arindam Roy, Shamik Sural, Arun Kumar Majumdar, Jaideep Vaidya, and Vijayalakshmi Atluri. Minimizing organizational user requirement while meeting security constraints. ACM Transactions on Management Information Systems (TMIS), 6(3):12:1–12:??, October 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Roy:2017:OEA

[RSM^{+}17] Arindam Roy, Shamik Sural, Arun Kumar Majumdar, Jaideep Vaidya, and Vijayalakshmi Atluri. On optimal employee assignment in constrained rolebased access control systems. ACM Transactions on Management Information Systems (TMIS), 7(4):10:1–10:??, January 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Roy:2021:OER

[RSM^{+}21] Arindam Roy, Shamik Sural, Arun Kumar Majumdar, Jaideep Vaidya, and Vijayalakshmi Atluri. Optimal employee recruitment in organizations under attribute-based access control. ACM Transactions on Management Information Systems (TMIS), 12(1):6:1–6:24, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3403950.

Rui:2011:DSB

[RW11] Huaxia Rui and Andrew Whinston. Designing a socialbroadcasting-based business intelligence system. ACM Transactions on Management Information Systems (TMIS), 2 (4):22:1–22:??, December 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Storey:2024:DSI

[SB24] Veda C. Storey and Richard Baskerville. Design with Simon’s inner and outer environments: Theoretical foundations for design science research methods for digital science. ACM Transactions on Management Information Systems (TMIS), 15(1):4:1–4:??, March 2024. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3640819.

Srinivasan:2023:HLS

[SCR23] Karthik Srinivasan, Faiz Currim, and Sudha Ram. A human-in-the-loop segmented mixed-effects modeling method for analyzing wearables data. ACM Transactions on Management Information Systems (TMIS), 14(2):18:1–18:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564276.

Sarker:2013:MOB

[SCT^{+}13] Suprateek Sarker, Suranjan Chakraborty, Patriya Silpakit Tansuhaj, Mark Mulder, and Kivilcim Dogerlioglu-Demir. The “Mail-Order-Bride” (MOB) phenomenon in the cyberworld: an interpretive investigation. ACM Transactions on Management Information Systems (TMIS), 4(3): 10:1–10:??, October 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Sudhakar:2021:DLM

[SG21] Tanuja Sudhakar and Marina Gavrilova. Deep learning for multi-instance biometric privacy. ACM Transactions on Management Information Systems (TMIS), 12(1):3:1–3:23, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3389683.

Siering:2023:PPP

[Sie23] Michael Siering. Peer-to-peer (P2P) lending risk management: Assessing credit risk on social lending platforms using textual factors. ACM Transactions on Management Information Systems (TMIS), 14(3): 25:1–25:??, 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3589003.

Srinivasan:2023:EDM

[SJ23] Karthik Srinivasan and Jinhang Jiang. Examining disease multimorbidity in U.S. hospital visits before and during COVID-19 pandemic: a graph analytics approach. ACM Transactions on Management Information Systems (TMIS), 14(2):17:1–17:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3564274.

Silva:2015:PAA

[SJC15] Thushari Silva, Ma Jian, and Yang Chen. Process analytics approach for R&D project selection. ACM Transactions on Management Information Systems (TMIS), 5(4):21:1–21:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Sun:2017:BCP

[SJKP17] Can Sun, Yonghua Ji, Bora Kolfal, and Ray Patterson. Business-to-consumer platform strategy: How vendor certification changes platform and seller incentives. ACM Transactions on Management Information Systems (TMIS), 8(2– 3):6:1–6:??, August 2017. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Shan:2012:OAC

[SK12] Zhe Shan and Akhil Kumar. Optimal adapter creation for process composition in synchronous vs. asynchronous communication. ACM Transactions on Management Information Systems (TMIS), 3(2): 8:1–8:??, July 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Samtani:2020:TAI

[SKC20] Sagar Samtani, Murat Kantarcioglu, and Hsinchun Chen. Trailblazing the artificial intelligence for cybersecurity discipline: a multi-disciplinary research roadmap. ACM Transactions on Management Information Systems (TMIS), 11 (4):17:1–17:19, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3430360.

Samtani:2021:MDP

[SKC21] Sagar Samtani, Murat Kantarcioglu, and Hsinchun Chen.

A multi-disciplinary perspective for conducting artificial intelligence-enabled privacy analytics: Connecting data, algorithms, and systems. ACM Transactions on Management Information Systems (TMIS), 12(1):1:1–1:18, March 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3447507.

Sutanto:2011:ESV

[SKT11] Juliana Sutanto, Atreyi Kankanhalli, and Bernard Cheng Yian Tan. Eliciting a sense of virtual community among knowledge contributors. ACM Transactions on Management Information Systems (TMIS), 2(3): 14:1–14:??, October 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Sutanto:2015:ITC

[SKT15] Juliana Sutanto, Atreyi Kankanhalli, and Bernard Cheng Yian Tan. Investigating task coordination in globally dispersed teams: a structural contingency perspective. ACM Transactions on Management Information Systems (TMIS), 6(2): 5:1–5:??, July 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Sweet:2020:VVC

[SMY20] Christopher Sweet, Stephen Moskal, and Shanchieh Jay Yang. On the variety and veracity of cyber intrusion alerts synthesized by generative adversarial networks. ACM Transactions on Management Information Systems (TMIS), 11 (4):22:1–22:21, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3394503.

Sainani:2020:IRS

[SNS^{+}20] Henanksha Sainani, Josephine M.

Namayanja, Guneeti Sharma, Vasundhara Misal, and Vandana P. Janeja. IP reputation scoring with geo-contextual feature augmentation. ACM Transactions on Management Information Systems (TMIS), 11(4):26:1–26:29, December 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3419373.

Sreenu:2020:CPP

[Sre20] Nenavath Sreenu. Cashless payment policy and its effects on economic growth of India: an exploratory study. ACM Transactions on Management Information Systems (TMIS), 11(3):15:1–15:10, August 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3391402.

Schmidt-Rauch:2011:TTA

[SRS11] Susanne Schmidt-Rauch and Gerhard Schwabe. From telesales to tele-advisory in travel agencies: Business problems, generic design goals and requirements. ACM Transactions on Management Information Systems (TMIS), 2(3): 17:1–17:??, October 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Sun:2016:UAN

[SSY16] Yutian Sun, Jianwen Su, and Jian Yang. Universal artifacts: a new approach to business process management (BPM) systems. ACM Transactions on Management Information Systems (TMIS), 7(1): 3:1–3:??, March 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Shao:2020:EEA

[STASH20] Sicong Shao, Cihan Tunc, Amany Al-Shawi, and Salim Hariri. An ensemble of ensembles approach to author attribution for Internet relay chat forensics. ACM Transactions on Management Information Systems (TMIS), 11(4):24:1– 24:25, December 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3409455.

Sun:2022:DAM

[STBI22] Bo Sun, Takeshi Takahashi, Tao Ban, and Daisuke Inoue. Detecting Android malware and classifying its families in large-scale datasets. ACM Transactions on Management Information Systems (TMIS), 13(2):12:1–12:21, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3464323.

Sutterer:2019:TBP

[SWB19] Paul Sutterer, Stefan Waldherr, and Martin Bichler. Are truthful bidders paying too much? Efficiency and revenue in display ad auctions. ACM Transactions on Management Information Systems (TMIS), 10(2):6:1–6:??, August 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3325523.

Sheng:2024:MMT

[SXH^{+}24] Jessica Qiuhua Sheng, Da Xu, Paul Jen-Hwa Hu, Liang Li, and Ting-Shuo Huang. Mining multimorbidity trajectories and co-medication effects from patient data to predict posthip fracture outcomes. ACM Transactions on Management Information Systems (TMIS), 15(2):10:1–10:??, June 2024. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3665250.

Sakata:2013:IEE

[SYS^{+}13] Masato Sakata, Zeynep Yücel, Kazuhiko Shinozawa, Norihiro Hagita, Michita Imai, Michiko Furutani, and Rumiko Matsuoka. An inference engine for estimating outside states of clinical test items. ACM Transactions on Management Information Systems (TMIS), 4(3): 13:1–13:??, October 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Shen:2023:ECA

[SZX^{+}23] Dazhong Shen, Hengshu Zhu, Keli Xiao, Xi Zhang, and Hui Xiong. Exploiting connections among personality, job position, and work behavior: Evidence from joint Bayesian learning. ACM Transactions on Management Information Systems (TMIS), 14(3):26:1–26:??, September 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3607875.

Tang:2022:QLW

[TCS22] Yan Tang, Weilong Cui, and Jianwen Su. A query language for workflow logs. ACM Transactions on Management Information Systems (TMIS), 13(2): 16:1–16:28, June 2022. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3482968.

Tan:2020:CPC

[TLK20] Liling Tan, Maggie Yundi Li, and Stanley Kok. E-commerce product categorization via machine translation. ACM Transactions on Management Information Systems (TMIS), 11(3): 11:1–11:14, August 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3382189.

Tsai:2014:SPS

[TQ14] Chih-Fong Tsai and Zen-Yu Quan. Stock prediction by searching for similarities in candlestick charts. ACM Transactions on Management Information Systems (TMIS), 5(2):9:1– 9:??, July 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Tuarob:2018:DDB

[TSCT18] Suppawong Tuarob, Ray Strong, Anca Chandra, and Conrad S. Tucker. Discovering discontinuity in big financial transaction data. ACM Transactions on Management Information Systems (TMIS), 9(1):3:1–3:??, February 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Tuzhilin:2011:KMR

[Tuz11] Alexander Tuzhilin. Knowledge management revisited: Old Dogs, New tricks. ACM Transactions on Management Information Systems (TMIS), 2(3): 13:1–13:??, October 2011. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Taghavi:2017:RCF

[TW17] Atefeh Taghavi and Carson Woo. The role clarity framework to improve requirements gathering. ACM Transactions on Management Information Systems (TMIS), 8(2–3): 9:1–9:??, August 2017. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Tsai:2016:DFK

[TWC16] Ming-Feng Tsai, Chuan-Ju Wang, and Po-Chuan Chien. Discovering finance keywords via continuous-space language models. ACM Transactions on Management Information Systems (TMIS), 7(3):7:1–7:??, October 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Tao:2020:WSW

[TZ20] Jie Tao and Lina Zhou. A weakly supervised WordNet-Guided deep learning approach to extracting aspect terms from online reviews. ACM Transactions on Management Information Systems (TMIS), 11(3): 13:1–13:22, August 2020. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3399630.

Teng:2021:ENF

[TZLX21] Mingfei Teng, Hengshu Zhu, Chuanren Liu, and Hui Xiong. Exploiting network fusion for organizational turnover prediction. ACM Transactions on Management Information Systems (TMIS), 12(2):16:1–16:18, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3439770.

Uhl:2011:EUC

[Uhl11] Matthias W. Uhl. Explaining U.S. consumer behavior with news sentiment. ACM Transactions on Management Information Systems (TMIS), 2(2): 9:1–9:??, June 2011. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Ullah:2013:SRB

[UL13] Azmat Ullah and Richard Lai.

A systematic review of business and information technology alignment. ACM Transactions on Management Information Systems (TMIS), 4(1):4:1– 4:??, April 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Unger:2023:DNA

[ULST23] Moshe Unger, Pan Li, Sahana (Shahana) Sen, and Alexander Tuzhilin. Don’t need all eggs in one basket: Reconstructing composite embeddings of customers from individual-domain embeddings. ACM Transactions on Management Information Systems (TMIS), 14(2):20:1–20:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578710.

Unger:2020:CAR

[UTL20] Moshe Unger, Alexander Tuzhilin, and Amit Livne. Contextaware recommendations based on deep learning frameworks. ACM Transactions on Management Information Systems (TMIS), 11(2):8:1–8:15, July 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3386243.

vanderAalst:2012:PMO

[vdA12]  Wil van der Aalst.  Pro-

cess mining: Overview and opportunities. ACM Transactions on Management Information Systems (TMIS), 3(2):7:1– 7:??, July 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

vanderLinden:2023:MVD

[vdLSFEA23] Sanne van der Linden, Rita Sevastjanova, Mathias Funk, and Mennatallah El-Assady. MediCoSpace: Visual decisionsupport for doctor-patient consultations using medical concept spaces from EHRs. ACM Transactions on Management Information Systems (TMIS), 14(2):15:1–15:??, June 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3564275.

Valecha:2013:DMC

[VSRU13] Rohit Valecha, Raj Sharman, H. Raghav Rao, and Shambhu Upadhyaya. A dispatch-mediated communication model for emergency response systems. ACM Transactions on Management Information Systems (TMIS), 4(1): 2:1–2:??, April 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Wang:2021:GCN

[WCL^{+}21] Xi Wang, Yibo Chai, Hui Li, Wenbin Wang, and Weishan Sun. Graph convolutional network-based model for incident-related congestion prediction: a case study of Shanghai expressways. ACM Transactions on Management Information Systems (TMIS), 12(3): 21:1–21:22, July 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3451356.

Wang:2021:ABB

[WCX21] Qin Wang, Shiping Chen, and Yang Xiang. Anonymous blockchain-based system for consortium. ACM Transactions on Management Information Systems (TMIS), 12(3):26:1– 26:25, July 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3459087.

Wang:2013:MTE

[WEM^{+}13] Zidong Wang, Julie Eatock, Sally McClean, Dongmei Liu, Xiaohui Liu, and Terry Young. Modeling throughput of emergency departments via time series: an expectation maximization algorithm. ACM Transactions on Management Information Systems (TMIS), 4 (4):16:1–16:??, December 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Wu:2013:DKM

[WH13] Jiming Wu and Clyde W. Holsapple. Does knowledge management matter? The empirical evidence from marketbased valuation. ACM Transactions on Management Information Systems (TMIS), 4(2): 6:1–6:??, August 2013. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Wen:2022:KFA

[WHEC22] Bo Wen, Paul Jen-Hwa Hu, Mohammadreza Ebrahimi, and Hsinchun Chen. Key factors affecting user adoption of openaccess data repositories in intelligence and security informatics: an affordance perspective. ACM Transactions on Management Information Systems (TMIS), 13(1):10:1–10:24, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460823.

Wei:2012:UNA

[WR12] Wei Wei and Sudha Ram. Using a network analysis approach for organizing social bookmarking tags and enabling Web content discovery. ACM Transactions on Management Information Systems (TMIS), 3(3):15:1–15:??, October 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Werder:2022:EDP

[WRZ22] Karl Werder, Balasubramaniam Ramesh, and Rongen (Sophia) Zhang. Establishing data provenance for responsible artificial intelligence systems. ACM Transactions on Management Information Systems (TMIS), 13(2):22:1– 22:23, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503488.

Wang:2015:AFU

[WWL^{+}15] G. Alan Wang, Harry Jiannan Wang, Jiexun Li, Alan S. Abrahams, and Weiguo Fan. An analytical framework for understanding knowledge-sharing processes in online Q&A communities. ACM Transactions on Management Information Systems (TMIS), 5(4):18:1– 18:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Wu:2022:OMS

[WWL^{+}22] Youxi Wu, Xiaohui Wang, Yan Li, Lei Guo, Zhao Li, Ji Zhang, and Xindong Wu. OWSP-Miner: Self-adaptive one-off weak-gap strong pattern mining. ACM Transactions on Management Information Systems (TMIS), 13(3):25:1–25:23, September 2022. CODEN ????

ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3476247.

Wu:2024:COO

[WWL^{+}24] Youxi Wu, Zhen Wang, Yan Li, Yingchun Guo, He Jiang, Xingquan Zhu, and Xindong Wu. Co-occurrence orderpreserving pattern mining with keypoint alignment for time series. ACM Transactions on Management Information Systems (TMIS), 15(2):9:1–9:??, June 2024. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3658450.

Wang:2010:DIS

[WXR10] Jingguo Wang, Nan Xiao, and H. Raghav Rao. Drivers of information security search behavior: an investigation of network attacks and vulnerability disclosures. ACM Transactions on Management Information Systems (TMIS), 1(1):3:1– 3:??, December 2010. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Wu:2022:EST

[WZW22] Xindong Wu, Xingquan Zhu, and Minghui Wu. The evolution of search: Three computing paradigms. ACM Transactions on Management Information Systems (TMIS), 13(2): 20:1–20:20, June 2022. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3495214.

Wang:2022:DDS

[WZZ22] Shui-Hua Wang, Xin Zhang, and Yu-Dong Zhang. DSSAE: Deep stacked sparse autoencoder analytical model for COVID-19 diagnosis by fractional Fourier entropy. ACM Transactions on Management Information Systems (TMIS), 13(1):2:1–2:20, March 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3451357.

Wang:2020:PUP

[WZZS20] Xiangyu Wang, Kang Zhao, Xun Zhou, and Nick Street. Predicting user posting activities in online health communities with deep learning. ACM Transactions on Management Information Systems (TMIS), 11(3):12:1–12:15, August 2020. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3383780.

Xiao:2018:PSD

[XLLX18] Keli Xiao, Qi Liu, Chuanren Liu, and Hui Xiong. Price shock detection with an influence-based model of social attention. ACM Transactions on Management Information Systems (TMIS), 9(1):2:1– 2:??, February 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Xu:2016:RMB

[XLZ^{+}16] Jiajie Xu, Chengfei Liu, Xiaohui Zhao, Sira Yongchareon, and Zhiming Ding. Resource management for business process scheduling in the presence of availability constraints. ACM Transactions on Management Information Systems (TMIS), 7(3):9:1–9:??, October 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Xue:2021:OOA

[XWC21] Xingsi Xue, Xiaojing Wu, and Junfeng Chen. Optimizing ontology alignment through an interactive compact genetic algorithm. ACM Transactions on Management Information Systems (TMIS), 12(2):14:1– 14:17, June 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3439772.

Xiong:2019:FES

[XWLC19] Hu Xiong, Yi Wang, Wenchao Li, and Chien-Ming Chen. Flexible, efficient, and secure access delegation in cloud computing. ACM Transactions on Management Information Systems (TMIS), 10 (1):2:1–2:??, May 2019. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3318212.

Xie:2021:WLP

[XZBZ21] Jiaheng Xie, Bin Zhang, Susan Brown, and Daniel Zeng. Write like a pro or an amateur? Effect of medical language formality. ACM Transactions on Management Information Systems (TMIS), 12(3):24:1– 24:25, July 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458752.

Xie:2022:RPP

[XZM^{+}22] Jiaheng Xie, Bin Zhang, Jian Ma, Daniel Zeng, and Jenny Lo-Ciganic. Readmission prediction for patients with heterogeneous medical history: a trajectory-based deep learning approach. ACM Transactions on Management Information Systems (TMIS), 13(2):14:1– 14:27, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468780.

Yaraghi:2013:NEH

[YDS^{+}13] Niam Yaraghi, Anna Ye Du, Raj Sharman, Ram D. Gopal, and R. Ramesh. Network effects in health information exchange growth. ACM Transactions on Management Information Systems (TMIS), 4(1):1:1– 1:??, April 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Yan:2015:MAG

[YHLW15] Jiaqi Yan, Daning Hu, Stephen S. Liao, and Huaiqing Wang. Mining agents’ goals in agentoriented business processes. ACM Transactions on Management Information Systems (TMIS), 5(4):20:1–20:??, January 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Yang:2013:SHW

[YLA13] Christopher C. Yang, Gondy Leroy, and Sophia Ananiadou. Smart health and wellbeing. ACM Transactions on Management Information Systems (TMIS), 4(4):15:1–15:??, December 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Yeo:2014:RMD

[YRUP14] M. Lisa Yeo, Erik Rolland, Jackie Rees Ulmer, and Raymond A. Patterson. Risk mitigation decisions for IT security. ACM Transactions on Management Information Systems (TMIS), 5(1):5:1–5:??, April 2014. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Yeo:2022:HCD

[YRUP22] M. Lisa Yeo, Erik Rolland, Jacquelyn Rees Ulmer, and Raymond A. Patterson. How customer demand reactions impact technology innovation and security. ACM Transactions on Management Information Systems (TMIS), 13(3):32:1–32:17, September 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505227.

Yang:2014:PDS

[YYJ14] Christopher C. Yang, Haodong Yang, and Ling Jiang. Postmarketing drug safety surveillance using publicly available health-consumer-contributed content in social media. ACM Transactions on Management Information Systems (TMIS), 5(1):2:1–2:??, April 2014. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Yu:2019:EAC

[YZJ^{+}19] Shuo Yu, Hongyi Zhu, Shan Jiang, Yong Zhang, Chunxiao Xing, and Hsinchun Chen. Emoticon analysis for Chinese social media and e-commerce: The AZEmo system. ACM Transactions on Management Information Systems (TMIS), 9(4):16:1–16:??, March 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3309707.

Zimbra:2018:SAT

[ZAZC18] David Zimbra, Ahmed Abbasi, Daniel Zeng, and Hsinchun Chen. The state-of-the-art in Twitter sentiment analysis: a review and benchmark evaluation. ACM Transactions on Management Information Systems (TMIS), 9(2):5:1– 5:??, September 2018. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zaeem:2021:EGP

[ZB21] Razieh Nokhbeh Zaeem and K. Suzanne Barber. The effect of the GDPR on privacy policies: Recent progress and future promise. ACM Transactions on Management Information Systems (TMIS), 12(1): 2:1–2:20, March 2021. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
doi/10.1145/3389685.

Zhu:2021:HRS

[ZBX^{+}21] Shixiang Zhu, Alexander Bukharin, Liyan Xie, Mauricio Santillana, Shihao Yang, and Yao Xie. High-resolution spatiotemporal model for countylevel COVID-19 activity in the U.S. ACM Transactions on Management Information Systems (TMIS), 12(4):33:1–33:20, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468876.

Zimbra:2015:SAF

[ZCL15] David Zimbra, Hsinchun Chen, and Robert F. Lusch. Stakeholder analyses of firm-related Web forums: Applications in stock return prediction. ACM Transactions on Management Information Systems (TMIS), 6(1):2:1–2:??, February 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zhang:2022:CDW

[ZELC22] Ning Zhang, Mohammadreza Ebrahimi, Weifeng Li, and Hsinchun Chen. Counteracting dark Web text-based CAPTCHA with generative adversarial learning for proactive cyber threat intelligence. ACM Transactions on Management Information Systems (TMIS), 13(2):21:1–21:21, June 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505226.

Zhang:2013:PCN

[ZGG13] Zhu Zhang, Chenhui Guo, and Paulo Goes. Product comparison networks for competitive analysis of online word-of-mouth. ACM Transactions on Management Information Systems (TMIS), 3(4):20:1–20:??, January 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zeng:2023:DAO

[ZJT^{+}23] Xiao Zeng, David Ji, Dimple R. Thadani, Boying Li, Xiaodie Pu, Zhao Cai, and Patrick Y. K. Chau. Disentangling affordances of online collaboration tools for mutual aid in emergencies: Insights from the COVID-19 lockdown. ACM Transactions on Management Information Systems (TMIS), 14(4):32:1–32:??, December 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3593056.

Zokaeinikoo:2021:AIA

[ZKMK21] Maryam Zokaeinikoo, Pooyan Kazemian, Prasenjit Mitra, and Soundar Kumara. AID-COV: an interpretable artificial intelligence model for detection of COVID-19 from chest radiography images. ACM Transactions on Management Information Systems (TMIS), 12 (4):29:1–29:20, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3466690.

Zhang:2012:DWM

[ZLC12] Zhu Zhang, Xin Li, and Yubo Chen. Deciphering word-of-mouth in social media: Textbased metrics of consumer reviews. ACM Transactions on Management Information Systems (TMIS), 3(1):5:1–5:??, April 2012. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zhao:2015:RBP

[ZLY^{+}15] Xiaohui Zhao, Chengfei Liu, Sira Yongchareon, Marek Kowalkiewicz, and Wasim Sadiq. Role-based process view derivation and composition. ACM Transactions on Management Information Systems (TMIS), 6(2):7:1–7:??, July 2015. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zhang:2013:MDP

[ZML^{+}13] He Zhang, Sanjay Mehotra, David Liebovitz, Carl A. Gunter, and Bradley Malin. Mining deviations from patient care pathways via electronic medical record system audits. ACM Transactions on Management Information Systems (TMIS), 4(4):17:1–17:??, December 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zo:2019:SOA

[ZNJ19] Hangjung Zo, Derek L. Nazareth, and Hemant K. Jain. Serviceoriented application composition with evolutionary heuristics and multiple criteria. ACM Transactions on Management Information Systems (TMIS), 10(3):10:1–10:??, November 2019. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3354288.

Zhang:2013:CMS

[ZTD^{+}13] Bin Zhang, Andrew C. Thomas, Patrick Doreian, David Krackhardt, and Ramayya Krishnan. Contrasting multiple social network autocorrelations for binary outcomes, with applications to technology adoption. ACM Transactions on Management Information Systems (TMIS), 3(4):18:1–18:??, January 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zhai:2023:RNB

[ZZ23] Shuang (Sophie) Zhai and Zhu (Drew) Zhang. Read the news, not the books: Forecasting firms’ long-term financial performance via deep text mining. ACM Transactions on Management Information Systems (TMIS), 14(1):3:1–3:??, March 2023. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3533018.

Zhang:2013:RWM

[ZZA^{+}13] Zhu Zhang, Daniel D. Zeng, Ahmed Abbasi, Jing Peng, and Xiaolong Zheng. A random walk model for item recommendation in social tagging systems. ACM Transactions on Management Information Systems (TMIS), 4(2):8:1–8:??, August 2013. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Zhu:2018:PJF

[ZZX^{+}18] Chen Zhu, Hengshu Zhu, Hui Xiong, Chao Ma, Fang Xie, Pengliang Ding, and Pan Li. Person-job fit: Adapting the right talent for the right job with joint representation learning. ACM Transactions on Management Information Systems (TMIS), 9(3):12:1– 12:??, November 2018. CO-DEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3234465.

Zhao:2021:ISS

[ZZYT21] Kang Zhao, Qingpeng Zhang, Sean H. Y. Yuan, and Kelvin Kam-Fai Tsoi. Introduction to the special section on using AI and data science to handle pandemics and related disruptions. ACM Transactions on Management Information Systems (TMIS), 12(4):27:1–27:2, December 2021. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486969.